### v3

# A functional central limit theorem for weighted occupancy processes of the Karlin model

Abstract.
A functional central limit theorem is established for weighted occupancy processes of the Karlin model. The weighted occupancy processes take the form of, with $D_{n,j}$ denoting the number of urns with $j$-balls after the first $n$ samplings, $\sum_{j=1}^{n}a_{j}D_{n,j}$ for a prescribed sequence of real numbers $(a_{j})_{j\in\mathbb{N}}$. The main applications are limit theorems for random permutations induced by Chinese restaurant processes with $(\alpha,\theta)$-seating with $\alpha\in(0,1),\theta>-\alpha$. An example is briefly mentioned here, and full details are provided in an accompanying paper.
1. Introduction
The infinite urn scheme has been a classical model in probability theory, dating back at least to the 1960s in the works of Bahadur, [3], Darling, [14] and Karlin, [27]. In particular, Darling and Karlin first considered frequencies decaying polynomially, that is, the sampling frequencies $(p_{j})_{j\in\mathbb{N}}$ of the infinite urn model satisfy $p_{j}=L(j)j^{-1/\alpha}$ for some slowly varying function $L$ at infinity and some parameter $\alpha\in(0,1]$. They also examined the case when $p_{j}$ decays faster than any polynomial rates (for a precise definition, see (1.2) below with $\alpha=0$). For these cases, Karlin established central limit theorems for various statistics, and therefore the model with such frequencies has been sometimes referred to as the Karlin model (although his seminal paper investigated other frequencies as well).
Among many statistics of interest, Karlin established central limit theorems for the so-called occupancy and odd-occupancy processes. If we let $D_{n,j}$ denote the number of urns with exactly $j$ balls after the first $n$ rounds of sampling, then these are the counting processes
(1.1) $S_{n}:=\sum_{j=1}^{n}D_{n,j}\quad\mbox{ and }\quad S_{n}^{\rm odd}:=\sum_{j=1}^{n}D_{n,j}{{\bf 1}}_{\left\{j\rm~{}is~{}odd\right\}},n\in\mathbb{N},$
respectively. See Gnedin et al., [20] for a survey on the infinite urn schemes. Surprisingly, the functional central limit theorems for the Karlin model have not been developed until the last 10 years, even Karlin’s results already indicated non-diffusive scaling limits for occupancy and odd-occupancy processes more than half a century ago. For functional central limit theorems on statistics of Karlin model and its variations, we mention [17, 12, 13, 16, 25, 26]. In particular, in [17] functional central limit theorems for occupancy and odd-occupancy processes have been established. Moreover, an additional randomization of the Karlin model was introduced and it was shown that the randomized odd-occupancy process scales to a fractional Brownian motion with Hurst index $H=\alpha/2\in(0,1/2)$. This result revealed a connection of the odd-occupancy processes to stochastic processes with long-range dependence [30, 32].
In the current paper, we establish a functional central limit theorem for the weighted sum of occupancy counts $(D_{n,j})_{j\in\mathbb{N}}$ with prescribed weights $\vec{a}=(a_{j})_{j\in\mathbb{N}}$ defined as
$S_{n}\equiv S_{n,\vec{a}}:=\sum_{j=1}^{n}a_{j}D_{n,j}.$
The statistics in this form include the (odd-)occupancy processes in (1.1) and hence we refer to $S_{n}$ as weighted occupancy processes. Throughout, $a_{0}=0$. In particular, with $a_{j}=1$ and $a_{j}={{\bf 1}}_{\left\{j~{}\rm~{}is~{}odd\right\}}$, $S_{n,\vec{a}}$ becomes the two cases in (1.1) considered by Karlin respectively; see also [17]. Another extensively studied case is $a_{j}={{\bf 1}}_{\left\{j\geq j_{0}\right\}}$, and $S_{n,\vec{a}}=\sum_{j\geq j_{0}}D_{n,j}$ is the counting process of all urns with at least $j_{0}$ balls. This case was again investigated already by Karlin; see also [12]. As an immediate consequence one obtains limit theorems for $D_{n,j_{0}}$ by simply writing $D_{n,j_{0}}=\sum_{j\geq j_{0}}D_{n,j}-\sum_{j\geq j_{0}+1}D_{n,j}$ and a continuous mapping argument. We are not aware of any other examples of $S_{n,\vec{a}}$ investigated in the literature.
For the general weights $(a_{j})_{j\in\mathbb{N}}$ we impose $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ for some $\beta\in[0,\alpha^{2}/2)$ and all $i,j\in\mathbb{N}\cup\{0\}$. When $\beta=0$ this implies that the weights are bounded, including the occupancy and odd-occupancy processes investigated earlier. Our main result is a functional central limit theorem for the counting process $(S_{\left\lfloor nt\right\rfloor})_{t\in[0,1]}$ appropriately normalized, and we identify a new family of Gaussian processes in the limit. Our results can be viewed as an extension of the functional central limit theorems in [17] (more precisely, Theorem 2.3 therein for the quenched version of the randomized model with fixed randomization $\varepsilon_{j}\equiv 1$, corresponding exactly the two cases in (1.1) here; see Remark 2.2 for more details). While we follow the standard Poissonization approach, it turned out that the analysis becomes much more involved even in the case where $(a_{j})_{j\in\mathbb{N}}$ is a bounded sequence. The technical challenges came from (a) proving the tightness for the Poissonized process where, due the fact that now $a_{j}$ may be unbounded (when $\beta>0$) we no longer have simple moment estimates, and (b) carrying out of the de-Possonization step, where we use a refined analysis on discrete local central limit theorems for Poisson and binomial distributions.
Our main motivation came from random matrix theory: using the general functional central limit theorem established here we shall investigate asymptotic behaviors of random permutation matrices induced by the Chinese restaurant processes with $(\alpha,\theta)$-seating. We briefly mention one application in Section 1.2: a functional central limit theorem for the linear statistics of eigenvalues of the aforementioned permutation matrices. In the accompanying paper [19] we provide full details and also another application on limit fluctuations of characteristic polynomials of the permutation matrices outside the support of its limit (which is the uniform measure on the unit circle $\{z\in\mathbb{C}:|z|=1\}$).
1.1. Main results
We first recall the setup of an infinite urn model with sampling frequencies $(p_{j})_{j\in\mathbb{N}}$, where $p_{1}\geq p_{2}\geq\cdots$ and $\sum_{j=1}^{\infty}p_{j}=1$. Each round, $Y_{i}$ is an independent sampling with $\mathbb{P}(Y_{i}=j)=p_{j}$ indicating the label of the urn that a ball is thrown in, and we consider the following statistics at each time $n\in\mathbb{N}$:
$K_{n,\ell}:=\sum_{i=1}^{n}{{\bf 1}}_{\left\{Y_{i}=\ell\right\}},\quad\ell\in\mathbb{N}\quad\mbox{ and }\quad D_{n,j}:=\sum_{\ell=1}^{\infty}{{\bf 1}}_{\left\{K_{n,\ell}=j\right\}},\quad j\in\mathbb{N}.$
In words, $K_{n,\ell}$ is the number of balls in the urn labeled by $\ell$ by round $n$ and $D_{n,j}$ is the number of urns with exactly $j$ balls by round $n$. The key here is the decay rate of the frequencies. Let $10<\alpha<$ 1. The following assumption was introduced first by Darling, [14, Section 4] and Karlin, [27] about the same time (Darling only investigated weak laws of large number and Karlin investigated central limit theorems)
(1.2) $v(x):=\sum_{i=1}^{\infty}\delta_{1/p_{i}}([0,x])=x^{\alpha}L(x),\quad x\geq 0,$
where $L$ is a slowly varying function at infinity. We note that if
(1.3) $p_{n}\sim\mathsf{C}_{0}n^{-1/\alpha},$
as $n\to\infty$ for some constant $\mathsf{C}_{0}>0$, then $v(x)\sim\mathsf{C}_{0}^{\alpha}x^{\alpha}$ as $x\rightarrow\infty$. Throughout we write $a_{n}\sim b_{n}$ if $\lim_{n\to\infty}a_{n}/b_{n}=1$. More generally, one can consider $p_{n}=n^{-1/\alpha}L(n)$ for some slowly varying function $L$ at infinity. However we do not pursue this generality, as for our main applications on random matrices, $L(n)$ converges to a constant almost surely, and hence (1.3) suffices. The proofs are already quite involved without considering a more general $L$.
We next introduce the limit process that will arise in our limit theorem. Let $\mathcal{M}_{\alpha}$ be a Gaussian random measure defined on +×$\mathbb{R}_{+}\times\Omega^{\prime}$ with intensity measure $\alpha r^{-\alpha-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$ and $(N^{\prime}(t))_{t\geq 0}$ be a standard Poisson process on $(\Omega^{\prime},\mathbb{P}^{\prime})$. Set
(1.4) $\mathcal{Z}_{\alpha}(t)\equiv\mathcal{Z}_{\alpha,\vec{a}}(t):=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left(a_{N^{\prime}(rt)}-{\mathbb{E}}^{\prime}a_{N^{\prime}(rt)}\right)\mathcal{M}_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),t\geq 0,$
for a certain sequence of real numbers $\vec{a}=(a_{j})_{j\in\mathbb{N}_{0}}$ with $\mathbb{N}_{0}=\{0,1,\dots\}$ and $a_{0}=0$. Note that the integrand can also be written as
$a_{N^{\prime}(rt)}-{\mathbb{E}}a_{N^{\prime}(rt)}=\sum_{j=1}^{\infty}a_{j}({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}^{\prime}(N^{\prime}(rt)=j)).$
We refer to [33] for background on stochastic integrals with respect to Gaussian random measures, and will comment on the conditions imposed on $(a_{j})_{j\in\mathbb{N}}$ in a moment. We only recall the facts that if $\mathcal{M}$ is a Gaussian random measure on a measurable space $(E,\mathcal{E})$ with control measure $\mu$, then $\int f{\rm d}\mathcal{M}$ is a centered Gaussian random variable for $f\in L^{2}(E,\mu)$ and ${\rm{Cov}}(\int f{\rm d}\mathcal{M},\int g{\rm d}\mathcal{M})=\int fg{\rm d}\mu$ with $f,g\in L^{2}(E,\mu)$.
For our functional central limit theorem, set
(1.5) $W_{n}(t):=\sum_{j=1}^{\left\lfloor nt\right\rfloor}a_{j}\left(D_{\left\lfloor nt\right\rfloor,j}-{\mathbb{E}}D_{\left\lfloor nt\right\rfloor,j}\right),t\in[0,1],n\in\mathbb{N}.$
Write also $\sigma_{n}:=v(n)^{1/2},n\in\mathbb{N}$. Under (1.3), we have $\sigma_{n}\sim\mathsf{C}_{0}^{\alpha/2}n^{\alpha/2}$. Throughout, the constant $C$ may change from line to line.
Theorem 1.1.
Assume that $(p_{j})_{j\in\mathbb{N}}$ satisfies (1.3) with $\alpha\in(0,1)$. Assume that there exists $\beta\in[0,\alpha^{2}/2)$ such that $(a_{j})_{j\in\mathbb{N}}$ satisfies that
(1.6) $|a_{i+j}-a_{i}|\leq Cj^{\beta},\quad\mbox{ for all }i,j\in\mathbb{N}_{0}.$
Then,
$\left(\frac{{W}_{n}(t)}{\sigma_{n}}\right)_{t\in[0,1]}\Rightarrow\left(\mathcal{Z}_{\alpha}(t)\right)_{t\in[0,1]},\mbox{ as }n\rightarrow\infty,$
in the space of càdlàg functions $D([0,1])$ with $J_{1}$ topology [7], where the process $\mathcal{Z}_{\alpha}$ is as in (1.4).
Remark 1.2.
We first comment on condition (1.6) on $(a_{j})_{j\in\mathbb{N}}$.
  1. (i)
    The condition (1.6) implies $|a_{j}|\leq Cj^{\beta}$, and this weaker condition alone with $\beta<\alpha/2$ is needed for $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ to be a family of well-defined Gaussian random variables. See Proposition 2.1. If say $a_{j}\sim Cj^{\beta}$ with $\beta>\alpha/2$ we expect a phase transition: a different normalization should be applied and a different stochastic process arises. This is left for future work.
  2. (ii)
    We impose condition (1.6) with $\beta<\alpha/2$ to show that the process $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ has a version with Hölder-continuous sample paths of index $\gamma<\alpha$. See again Proposition 2.1.
  3. (iii)
    Condition (1.6) with the more restrictive constraint $\beta<\alpha^{2}/2$ (recall $\alpha\in(0,1)$) is needed to show that the Poissonized model is tight in $D([0,1])$, which is necessary to show the functional central limit theorem for the Poissonized version in Section 3.2. Assuming the weaker condition $|a_{j}|\leq Cj^{\beta}$ with $\beta<\alpha/2$ and no condition on the increments, we managed to prove the convergence of finite-dimensional distributions for the Poissonized model in Proposition 3.2; although this result alone is not enough to prove Theorem 1.1 (in fact, even the central limit theorem with $t=1$ fixed cannot be obtained). Note that we also need (1.6) with $\beta<\alpha/2$ in the de-Poissonization step. It is plausible that the constraint $\beta<\alpha^{2}/2$ can be relaxed to $\beta<\alpha/2$ in Section 3.2.
  4. (iv)
    It is not clear to us when $a_{j}$ is unbounded whether the functional central limit theorem can be established without any further conditions beside $|a_{j}|\leq Cj^{\beta}$. It is known that when $a_{j}$ changes sign, the regularity of the sample paths can be quite poor. For a closely related issue, see [16, Remark 2].
Remark 1.3.
For each $j\in\mathbb{N}$, set
$\mathcal{Z}_{\alpha,j}(t):=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}^{\prime}(N^{\prime}(rt)=j))\mathcal{M}_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),t\geq 0.$
Our process $\mathcal{Z}_{\alpha}$ is in fact a linear combination of the above with coefficients $(a_{j})_{j\in\mathbb{N}}$. That is,
$\sum_{j=1}^{\infty}a_{j}\mathcal{Z}_{\alpha,j}(t)=\int_{\mathbb{R}_{+}\times\Omega_{+}}\sum_{j=1}^{\infty}a_{j}({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}^{\prime}(N^{\prime}(rt)=j))\mathcal{M}_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),t\geq 0,$
where the left-hand side is understood as the $L^{2}$-limit of the infinite series and the right-hand side is a stochastic integral.
The processes $\mathcal{Z}_{\alpha,j}$ and their linear combinations have appeared in earlier investigations of Karlin models. It is worth noting in the literature, limiting Gaussian processes are often characterized by their explicit covariance functions, with no stochastic-integral representations provided. However, stochastic-integral representations offer additional insight into Gaussian processes. In fact, in [17], the stochastic-integral representations can only be read implicitly in the proofs, and they are explained explicitly only later in [18].
Remark 1.4.
We mention the recent paper by Iksanov and Kotelnikova, [26] (see also [25, Proposition 4.1]), who investigated Karlin model with $\alpha=0$ in (1.2) (note that in this case (1.3) does not make sense; but instead $p_{j}$ decays faster than any polynomial rates), and obtained new Gaussian processes as the scaling limits for various statistics. In short, the case $\alpha=0$ of the Karlin model behaves drastically different from the case $\alpha\in(0,1)$ investigated here, yet it appears to have received much less attention. The Karlin model with $\alpha=0$ is closely related to the Bernoulli sieves [21, 24]. See also [26, 25] for related references.
1.2. An application in random matrix theory
The key connection to random permutation matrices is that, thanks to Kingman’s representation theorem [28], the occupancy counts $(D_{n,1},\dots,D_{n,n})$ of an infinite urn scheme have the same joint law of the cycle counts $(C_{n,1},\dots,C_{n,n})$ of the corresponding exchangeable random partition from a Chinese restaurant process (where $C_{n,j}$ is the number of the $j$-cycles of the partition of $[n]=\{1,2,...,n\}$), if the rate $(p_{j})_{j\in\mathbb{N}}$ is appropriately chosen (strictly speaking, this statement involves regular conditional probability; see [19] for more details). Since for a permutation the cycle structure determines completely the eigenvalues of the corresponding matrix, many statistics of interest can be expressed as weighted sum of cycle counts, which have the same law as the weighted occupancy counts investigated here. This connection allows us to study random permutation matrices by investigating classical questions on infinite urn schemes. This idea was first due to Diaconis and Shahshahani, [15] and has been applied extensively to random permutation matrices either uniform or those induced by Ewens measures on symmetric groups with parameter $\theta>0$, corresponding to the Chinese restaurant processes with $(0,\theta)$-seating (see e.g. [1, 4, 6, 35, 5]). At the same time, the Karlin model with parameter $\alpha\in(0,1)$ can be associated to Chinese restaurant processes with $(\alpha,\theta)$-seating ($\theta>-\alpha$), and this indicates again the importance of Karlin model from another aspect. Surprisingly, we have found very few results on random permutations induced by $(\alpha,\theta)$-seating with $\alpha\in(0,1)$, and the connection to the infinite urn model has not been fully exploited yet.
Here we mention briefly an application in our accompanying paper [19], just to provide a flavor of the results to be established. We skip most of the background including the definition of Chinese restaurant processes, as more details and other applications can be found therein.
A Chinese restaurant process consists of a sequence of exchangeable random partitions, $(\Pi_{n})_{n\in\mathbb{N}}$, each is a random partition of $[n]$. Assuming further a uniformly random ordering of each block of the partition leads to a random permutation of $[n]$, we let $M_{n}$ denote the associated permutation matrix. Our goal is to study the asymptotic behavior of eigenvalues of $M_{n}$ as $n\to\infty$. Namely for each $M_{n}$, letting $e^{{\rm i}2\pi\theta_{n,1}},\dots,e^{{\rm i}2\pi\theta_{n,n}}$ represent its $n$ eigenvalues, with $\theta_{n,j}\in[0,1),j=1,\dots,n$, we are interested in the asymptotic behavior of the following linear statistics of the spectrum of a matrix $M_{n}$ for some test function $f:[0,1)\to\mathbb{R}$:
$S_{n}(f):=\sum_{i=1}^{n}f(\theta_{n,i}).$
Before stating the limit theorem, a key notion that needs to be recalled is the asymptotic frequencies of the random partitions $(\Pi_{n})_{n\in\mathbb{N}}$, which, ordered in decreasing order are denoted by $(P_{j}^{\downarrow})_{j\in\mathbb{N}}$. (In words, $P_{j}$ without ordering is the asymptotic proportion of customers sitting at the $j$-th table.) It is well-known that they are non-deterministic, strictly positive, and satisfy $\sum_{i=1}^{\infty}P_{i}^{\downarrow}=1$, and in particular for $\alpha\in(0,1),\theta>-\alpha$,
$P_{n}^{\downarrow}\sim\left(\frac{\mathsf{s}_{\alpha,\theta}}{\Gamma(1-\alpha)}\right)^{1/\alpha}n^{-1/\alpha},\quad\mbox{ almost surely.}$.
Here and throughout, $\Gamma$ is the gamma function. The random variable $\mathsf{s}_{\alpha,\theta}$ referred to as the $\alpha$-diversity, is strictly positive and is measurable with respect to
${\mathcal{P}}:=\sigma\left((P_{i}^{\downarrow})_{i\in\mathbb{N}}\right).$
The limit theorems on random matrices are stated in the notion of almost sure weak convergence. Given a Polish space $(\mathcal{X},d)$ and random elements $(X_{n})_{n\in\mathbb{N}},X$ in it, we say $X_{n}$ converges almost surely weakly to $X$ as $n\to\infty$ with respect to ${\mathcal{P}}$, denoted by $X_{n}\stackrel{{\scriptstyle a.s.w.}}{{\to}}X$ with respect to ${\mathcal{P}}$, if for all continuous and bounded functions $g$ on $\mathcal{X}$, $\lim_{n\to\infty}{\mathbb{E}}\left(g(X_{n})\mid{\mathcal{P}}\right)={\mathbb{E}}\left(g(X)\mid{\mathcal{P}}\right)$ almost surely. For a slightly more general notion, see [22].
Introduce
$a_{j}(f):=jR_{j}(f)\quad\mbox{ with }\quad R_{j}(f):=\frac{1}{j}\sum_{k=0}^{j-1}f\left(\frac{k}{j}\right)-\int_{0}^{1}f(x){\rm d}x,j\in\mathbb{N}.$
Theorem 1.5.
For functions $f:[0,1)\to\mathbb{R}$ such that $(a_{j})_{j\in\mathbb{N}}\equiv(a_{j}(f))_{j\in\mathbb{N}}$ satisfy the assumptions in Theorem 1.1, we have
$\left(\frac{S_{\left\lfloor nt\right\rfloor}(f)-{\mathbb{E}}\left(S_{\left\lfloor nt\right\rfloor}(f)\mid{\mathcal{P}}\right)}{n^{\alpha/2}}\right)_{t\in[0,1]}\stackrel{{\scriptstyle a.s.w.}}{{\to}}\left(\frac{\mathsf{s}_{\alpha,\theta}}{\Gamma(1-\alpha)}\right)^{1/2}\left(\mathcal{Z}_{\alpha}(t)\right)_{t\in[0,1]}$
with respect to ${\mathcal{P}}$ in $D([0,1])$, where on the right-hand side the process $\mathcal{Z}_{\alpha}$ is as in (1.4) and independent from ${\mathcal{P}}$.
The class of functions satisfying $|a_{j}(f)|\leq C$ include for example indicator functions $f={\bf 1}_{[a,b)}$ and the class $W^{1,p}([0,1])$ with $p\in[1,\infty]$ (absolutely continuous functions with $f^{\prime}\in L^{p}([0,1])$).
Our results are to be compared with limit theorems for linear statistics of random permutation matrices following Ewens measures by Ben Arous and Dang, [6]. There are two noticeable differences. First, they assumed $\sum_{j=1}^{\infty}a_{j}^{2}/j=\infty$ and $\max_{j=1,\dots,n}|a_{j}|=o(s_{n})$ with $s_{n}=(\sum_{j=1}^{n}a_{j}^{2}/j)^{1/2}$ (guaranteed by the assumption that $f$ is of bounded variation), and in this case it was shown that the normalization for the central limit theorem is $s_{n}$ which is of order at most $O(\sqrt{\log n})$ (this is the Gaussian regime, and when $\sum_{j=1}^{\infty}a_{j}^{2}/j<\infty$ the limit is a weighted sum of centered Poisson random variables). Here, the normalization is $n^{\alpha/2}$. Second, they only established a central limit theorem (for $t=1$ fixed), and their method based on Feller’s coupling does not allow one to establish a functional central limit theorem as ours, which provide additional information on the temporal evolution of eigenvalues. (Instead, the Feller’s coupling allows one to establish another type of functional central limit theorem as in Hansen, [23].)
The rest of the paper is organized as follows. Section 2 provides more properties on the limit process $\mathcal{Z}$. Section 3 provides the proof for the Poissonized model. Section 4 provides the proof of the de-Poissonization step.
Acknowledgements
The authors would like to thank two anonymous referees for very careful reading of and helpful comments on the manuscript. Y.W. was grateful to Alexander Iksanov for several illuminating discussions. J.G. and Y.W. were partially supported by Army Research Office, US (W911NF-20-1-0139). Y.W. was also partially supported by Simons Foundation (MP-TSM-00002359), and a Taft Center Fellowship (2024–2025) from Taft Research Center at University of Cincinnati.
2. The limit process
Recall the process $\mathcal{Z}_{\alpha}=\mathcal{Z}_{\alpha,\vec{a}}$ defined in (1.4), and that it depends on a sequence $(a_{j})_{j\in\mathbb{N}}$ with $a_{0}=0$.
Proposition 2.1.
Assume
(2.1) $|a_{j}|\leq Cj^{\beta}\quad\mbox{ with }\quad\beta\in\Big{[}0,\frac{\alpha}{2}\Big{)},$
Then, $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ is a family of centered Gaussian random variables with
(2.2) ${\rm{Cov}}(\mathcal{Z}_{\alpha}(s),\mathcal{Z}_{\alpha}(t))=\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\Gamma(k-\alpha)}{k!}-\alpha(s+t)^{\alpha}\sum_{k=2}^{\infty}\frac{F_{\vec{a},s,t}^{(2)}(k)\Gamma(k-\alpha)}{k!},$
where $0<s<t$, and
$F_{\vec{a},s,t}^{(1)}(k)$ $:=\sum_{j=1}^{k}\binom{k}{j}\left(\frac{s}{t}\right)^{j}\left(1-\frac{s}{t}\right)^{k-j}a_{j}a_{k},$
$F_{\vec{a},s,t}^{(2)}(k)$ $:=\sum_{j=1}^{k-1}\binom{k}{j}\left(\frac{s}{s+t}\right)^{j}\left(\frac{t}{s+t}\right)^{k-j}a_{j}a_{k-j}.$
If furthermore
(2.3) $|a_{i+j}-a_{i}|\leq Cj^{\beta}\mbox{ for all }i,j\in\mathbb{N}\quad\mbox{ with}\quad\beta\in\Big{[}0,\frac{\alpha}{2}\Big{)},$
then $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ forms a stochastic process with a version that is $\gamma$-Hölder-continuous for every $\gamma\in(0,\alpha/2)$.
Note that the assumption (2.3) includes the case that $(a_{j})_{j\in\mathbb{N}}$ is a sub-additive sequence with polynomial growing rate at most $\beta$.
Proof.
We compute directly the covariance, and will see at the end that it is a finite value under the assumption (2.1). By property of stochastic integrals,
${\rm{Cov}}(\mathcal{Z}_{\alpha}(s),\mathcal{Z}_{\alpha}(t))$ $=\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left(\left(a_{N^{\prime}(rs)}-{\mathbb{E}}^{\prime}a_{N^{\prime}(rs)}\right)\left(a_{N^{\prime}(rt)}-{\mathbb{E}}^{\prime}a_{N^{\prime}(rt)}\right)\right)\alpha r^{-\alpha-1}{\rm d}r$
(2.4) $=\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left(a_{N^{\prime}(rs)}a_{N^{\prime}(rt)}\right)\alpha r^{-\alpha-1}{\rm d}r-\int_{0}^{\infty}{\mathbb{E}}^{\prime}a_{N^{\prime}(rs)}{\mathbb{E}}^{\prime}a_{N^{\prime}(rt)}\alpha r^{-\alpha-1}{\rm d}r.$
Here and below, ${\mathbb{E}}^{\prime}$ is the expectation with respect to $\mathbb{P}^{\prime}$. The notation is solely to indicate that they have nothing to do with the probability space where the random measure $\mathcal{M}_{\alpha}$ is defined and for convenience from now on, when working with ${\mathbb{E}}^{\prime},\mathbb{P}^{\prime}$ and $N^{\prime}$ without the involvement of $\mathcal{M}_{\alpha}$ we may simply write ${\mathbb{E}},\mathbb{P}$ and $N$ accordingly. Recall that $N^{\prime}(rs)\leq N^{\prime}(rt)$ a.s. since $s<t$. The first integral above equals
$\int_{0}^{\infty}$ $\sum_{j=1}^{\infty}\sum_{j^{\prime}=j}^{\infty}a_{j}a_{j^{\prime}}\frac{(rs)^{j}}{j!}e^{-rs}\frac{(r(t-s))^{j^{\prime}-j}}{(j^{\prime}-j)!}e^{-r(t-s)}\alpha r^{-\alpha-1}{\rm d}r$
$=\alpha\sum_{k=1}^{\infty}\sum_{j=1}^{k}\binom{k}{j}a_{j}a_{k}s^{j}(t-s)^{k-j}\frac{1}{k!}\int_{0}^{\infty}e^{-rt}r^{k-\alpha-1}{\rm d}r$
$=\alpha t^{\alpha}\sum_{k=1}^{\infty}\sum_{j=1}^{k}\binom{k}{j}a_{j}a_{k}\left(\frac{s}{t}\right)^{j}\left(1-\frac{s}{t}\right)^{k-j}\frac{\Gamma(k-\alpha)}{k!}=\alpha t^{\alpha}\sum_{k=1}^{\infty}F_{\vec{a},s,t}^{(1)}(k)\frac{\Gamma(k-\alpha)}{k!}.$
Note that in the last step, we used the fact that $|F_{\vec{a},s,t}^{(1)}(k)|\leq Ck^{2\beta}$ under the assumption (2.1); it then follows that the last series is absolutely convergent (recall that $\Gamma(k-\alpha)/k!\sim k^{-1-\alpha}$ as $k\to\infty$), and hence the interchange of integral and summations in previous steps is also justified by Fubini’s theorem.
Similarly, the second integral in (2.4) equals
$\int_{0}^{\infty}$ $\sum_{j=1}^{\infty}\sum_{j^{\prime}=1}^{\infty}a_{j}a_{j^{\prime}}\frac{(rs)^{j}(rt)^{j^{\prime}}}{j!j^{\prime}!}e^{-r(s+t)}\alpha r^{-\alpha-1}{\rm d}r$
$=\alpha\sum_{k=2}^{\infty}\sum_{j=1}^{k-1}\binom{k}{j}a_{j}a_{k-j}s^{j}t^{k-j}\frac{1}{k!}\int_{0}^{\infty}e^{-r(s+t)}r^{k-\alpha-1}{\rm d}r$
$=\alpha(s+t)^{\alpha}\sum_{k=2}^{\infty}\sum_{j=1}^{k-1}\binom{k}{j}a_{j}a_{k-j}\left(\frac{s}{s+t}\right)^{j}\left(\frac{t}{s+t}\right)^{k-j}\frac{\Gamma(k-\alpha)}{k!}$
$=\alpha(s+t)^{\alpha}\sum_{k=2}^{\infty}F_{\vec{a},s,t}^{(2)}(k)\frac{\Gamma(k-\alpha)}{k!},$
and we used the fact that $|F_{\vec{a},s,t}^{(2)}(k)|\leq Ck^{2\beta}$ under the assumption (2.1). We have thus proved (2.2).
Now we examine the path continuity. By a similar calculation leading to (2.4), we have
${\mathbb{E}}\left(\mathcal{Z}_{\alpha}(t)-\mathcal{Z}_{\alpha}(s)\right)^{2}$ $=\int_{0}^{\infty}{\mathbb{E}}\left(a_{N(rt)}-a_{N(rs)}\right)^{2}\alpha r^{-\alpha-1}{\rm d}r$
$\quad-\int_{0}^{\infty}\left({\mathbb{E}}a_{N(rt)}-{\mathbb{E}}a_{N(rs)}\right)^{2}\alpha r^{-\alpha-1}{\rm d}r$
$\leq\int_{0}^{\infty}{\mathbb{E}}\left(a_{N(rt)}-a_{N(rs)}\right)^{2}\alpha r^{-\alpha-1}{\rm d}r.$
By (2.3), it follows that, for all $0<s<t$,
${\mathbb{E}}\left(\mathcal{Z}_{\alpha}(t)-\mathcal{Z}_{\alpha}(s)\right)^{2}$ $\leq C\int_{0}^{\infty}{\mathbb{E}}(N(r(t-s)))^{2\beta}\alpha r^{-\alpha-1}{\rm d}r$
$=C\int_{0}^{\infty}\sum_{k=1}^{\infty}k^{2\beta}\frac{(r(t-s))^{k}}{k!}e^{-r(t-s)}\alpha r^{-\alpha-1}{\rm d}r$
$=C\alpha(t-s)^{\alpha}\sum_{k=1}^{\infty}\frac{k^{2\beta}\Gamma(k-\alpha)}{\Gamma(k+1)}=C(t-s)^{\alpha}.$
Also, by well-known relation of moments for Gaussian distributions, we have
${\mathbb{E}}\left(\mathcal{Z}_{\alpha}(t)-\mathcal{Z}_{\alpha}(s)\right)^{2m}\leq C_{m}\left({\mathbb{E}}\left(\mathcal{Z}_{\alpha}(t)-\mathcal{Z}_{\alpha}(s)\right)^{2}\right)^{m}\leq C(t-s)^{\alpha m}.$
Then for $m>1/\alpha$, Kolmogorov’s continuity theorem entails that the process has a continuous version with $\gamma$-Hölder-continuous sample path with $\gamma<(\alpha m-1)/(2m)$. Letting $m\to\infty$, we complete the proof for the second part. ∎
Remark 2.2.
From stochastic-integral representations, we obtained an explicit formula of the covariance function (2.2). However, most of the time the infinite series in the formulae cannot be simplified and are not directly useful. We mention two noteworthy exceptions, both from [17, Theorem 2.3].
  1. (a)
    With $a_{j}=1,j\in\mathbb{N}$, $\mathcal{Z}_{\alpha}$ is the limit of occupancy process. Then,
    (2.5) ${\rm{Cov}}(\mathcal{Z}_{\alpha}(s),\mathcal{Z}_{\alpha}(t))=\Gamma(1-\alpha)\left((s+t)^{\alpha}-\max\{s,t\}^{\alpha}\right).$
  2. (b)
    With $a_{j}={{\bf 1}}_{\left\{j\rm~{}is~{}odd\right\}}$, $\mathcal{Z}_{\alpha}$ is the limit of odd-occupancy process. Then,
    ${\rm{Cov}}(\mathcal{Z}_{\alpha}(s),\mathcal{Z}_{\alpha}(t))=\Gamma(1-\alpha)2^{\alpha-2}\left((s+t)^{\alpha}-|t-s|^{\alpha}\right).$
    In this case, $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ is known as the bi-fractional Brownian motion with parameters $(1/2,\alpha)$. See [17, Remark 2.4] for more discussions.
We provide a calculation for (2.5) using (2.2). Assume $s<t$. First,
$F_{\vec{a},s,t}^{(1)}(k)$ $=\sum_{j=1}^{k}\binom{k}{j}\left(\frac{s}{t}\right)^{j}\left(1-\frac{s}{t}\right)^{k-j}=1-\left(1-\frac{s}{t}\right)^{k},$
$F_{\vec{a},s,t}^{(2)}(k)$ $=\sum_{j=1}^{k-1}\binom{k}{j}\left(\frac{s}{s+t}\right)^{j}\left(\frac{t}{s+t}\right)^{k-j}=1-\left(\frac{t}{s+t}\right)^{k}-\left(\frac{s}{s+t}\right)^{k}.$
A key identity in calculations involving limit processes of Karlin model is the following. Let $Q_{\alpha}$ be an $\alpha$-Sibuya random variable [31]. That is, it is an $\mathbb{N}$-valued random variable with
$\mathbb{P}(Q_{\alpha}=k)=\frac{\alpha}{\Gamma(1-\alpha)}\frac{\Gamma(k-\alpha)}{\Gamma(k+1)},k\in\mathbb{N},$
and in particular it satisfies $z^{\alpha}=1-{\mathbb{E}}(1-z)^{Q_{\alpha}}$ for all $z\in[0,1)$ (this is essentially Newton’s generalized binomial theorem). Then, we have
$\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\Gamma(k-\alpha)}{\Gamma(k+1)}=\Gamma(1-\alpha)t^{\alpha}{\mathbb{E}}\left(1-\left(1-\frac{s}{t}\right)^{Q_{\alpha}}\right)=\Gamma(1-\alpha)s^{\alpha},$
and
$\alpha(s+t)^{\alpha}\sum_{k=2}^{\infty}\frac{F_{\vec{a},s,t}^{(2)}(k)\Gamma(k-\alpha)}{\Gamma(k+1)}$ $=\Gamma(1-\alpha)(s+t)^{\alpha}{\mathbb{E}}\left(1-\left(\frac{t}{s+t}\right)^{Q_{\alpha}}-\left(\frac{s}{s+t}\right)^{Q_{\alpha}}\right)$
$=\Gamma(1-\alpha)\left(s^{\alpha}+t^{\alpha}-(s+t)^{\alpha}\right).$
Combining the above we have derived the stated (2.5).
3. Limit theorem for the Poissonized model
In order to prove Theorem 1.1, we proceed by a standard Poissonization approach, as already carried out in [17, 16, 13]. This section is devoted to the functional central limit theorem for the Poissonized model of the Karlin model, which we first introduce.
For each $\ell\in\mathbb{N}$, let $N_{\ell}=(N_{\ell}(t))_{t\geq 0}$ be a Poisson process with parameter $p_{\ell}$, and assume that $(N_{\ell})_{\ell\in\mathbb{N}}$ are independent. Set
$\widetilde{D}_{t,j}:=\sum_{\ell=1}^{\infty}{{\bf 1}}_{\left\{N_{\ell}(t)=j\right\}},t\geq 0,j\in\mathbb{N},$
and define
(3.1) $\widetilde{W}_{n}(t):=\sum_{j=1}^{\infty}a_{j}\left(\widetilde{D}_{nt,j}-\mathbb{E}\widetilde{D}_{nt,j}\right)=\sum_{\ell=1}^{\infty}\left(a_{N_{\ell}(nt)}-\mathbb{E}a_{N_{\ell}(nt)}\right).$
We shall prove the following theorem, which is the Poissonized version of Theorem 1.1. Recall that we assume $p_{j}\sim\mathsf{C}_{0}j^{-1/\alpha}$ as $j\to\infty$, which we do not repeat in the sequel, and we shall denote $\sigma_{n}^{2}:=v(n)\sim\mathsf{C}_{0}^{\alpha}n^{\alpha}$. See the discussions around (1.3).
Theorem 3.1.
Assume $(a_{j})_{j\in\mathbb{N}}$ with $a_{0}=0$ satisfies $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ for all $i,j\in\mathbb{N}$ and some $\beta\in[0,\alpha^{2}/2)$. Then,
$\left(\frac{\widetilde{W}_{n}(t)}{\sigma_{n}}\right)_{t\in[0,1]}\Rightarrow\left(\mathcal{Z}_{\alpha}(t)\right)_{t\in[0,1]},\mbox{ as }n\rightarrow\infty,$
in $D([0,1])$, where the process $\mathcal{Z}_{\alpha}$ is as in Proposition 2.1.
To establish Theorem 3.1, we first prove the convergence of finite-dimensional distributions in Section 3.1. Note that in this step we assume only the weaker condition $|a_{j}|\leq Cj^{\beta}$ for $\beta\in[0,\alpha/2)$. Next, in Section 3.2 we establish the tightness, and it is here that we impose the stronger condition $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ with $\beta\in[0,\alpha^{2}/2)$.
3.1. Finite-dimensional convergence for the Poissonized model
In this section, we prove the following.
Proposition 3.2.
Suppose that $(a_{j})_{j\in\mathbb{N}}$ satisfies $a_{j}\leq Cj^{\beta}$ for some $\beta\in[0,\alpha/2)$. Then,
$\left(\frac{\widetilde{W}_{n}(t)}{\sigma_{n}}\right)_{t\in[0,1]}\xrightarrow[]{f.d.d}\left(\mathcal{Z}_{\alpha}(t)\right)_{t\in[0,1]},\text{ as }n\rightarrow\infty.$
We proceed by the following approximation. Throughout this subsection without loss of generality assume that $(p_{\ell})_{\ell\in\mathbb{N}}$ is non-increasing. For every fixed $\epsilon>0$, write
$\widetilde{W}_{n}(t)=\widetilde{W}^{(\epsilon)}_{n}(t)+\widetilde{T}^{(\epsilon)}_{n}(t),\quad t\in[0,1],n\in\mathbb{N},$
where
$\widetilde{W}^{(\epsilon)}_{n}(t):=\sum_{\ell=\ell_{n}+1}^{\infty}\left({a}_{N_{\ell}(nt)}-{\mathbb{E}}a_{N_{\ell}(nt)}\right)\quad\mbox{ and }\quad\widetilde{T}^{(\epsilon)}_{n}(t):=\sum_{\ell=1}^{\ell_{n}}\left({a}_{N_{\ell}(nt)}-{\mathbb{E}}a_{N_{\ell}(nt)}\right),$
with
$\ell_{n}\equiv\ell_{n,\epsilon}:=\left\lfloor\epsilon v(n)\right\rfloor=\left\lfloor\epsilon\sigma_{n}^{2}\right\rfloor.$
Introduce also
$\mathcal{Z}_{\alpha}^{(\epsilon)}(t):=\int_{(0,\epsilon^{-1/\alpha}]\times\Omega^{\prime}}\left(a_{N^{\prime}(rt)}-{\mathbb{E}}^{\prime}a_{N^{\prime}(rt)}\right)\mathcal{M}_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),\quad j\in\mathbb{N},t\in[0,1].$
Clearly, for every $t\geq 0$ fixed, $\mathcal{Z}_{\alpha}^{(\epsilon)}(t)\to\mathcal{Z}_{\alpha}(t)$ in probability as $\epsilon\downarrow 0$.
Then, by the standard triangular array approximation argument [7, Theorem 3.1], to prove Proposition 3.2 it suffices to establish Lemmas 3.3 and 3.4 below.
Lemma 3.3.
For every $t>0$, under the assumptions of Proposition 3.2 we have
[…]
Lemma 3.4.
Under the assumptions of Proposition 3.2, for every $t>0$,
$\lim_{\epsilon\downarrow 0}\limsup_{n\to\infty}\mathbb{P}\left(\left|\widetilde{T}_{n}^{(\epsilon)}(t)\right|>\sigma_{n}\eta\right)=0,\mbox{ for all }\eta>0.$
Before proving the lemmas we need the following technical estimate which we shall use at several places. Let
$\gamma(a,t)=\int_{0}^{t}x^{a-1}e^{-x}{\rm d}x,\quad a>0,t>0,$
denote the lower incomplete gamma function.
Lemma 3.5.
Let $(b_{j})_{j\in\mathbb{N}_{0}}$ be a sequence such that $b_{0}=0$, $|b_{j}|\leq Cj^{\alpha^{\prime}}$ for some $\alpha^{\prime}<\alpha$. Fix $t>0$. Then,
(3.2) $\sum_{\ell=\ell_{n}+1}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}\sim\sigma_{n}^{2}\alpha t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\gamma(j-\alpha,t\epsilon^{-{1}/{\alpha}})}{j!},$
as $n\to\infty$, and there exists a constant $C>0$ such that
(3.3) $\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}|b_{N_{\ell}(nt)}|\leq C\sigma_{n}^{2}t^{\alpha^{\prime}}\epsilon^{1-\alpha^{\prime}/\alpha}$
for all $n\in\mathbb{N}$. As a consequence, there exists a constant $C>0$ such that for all $t\in(0,1]$,
$\sum_{\ell=1}^{\infty}{\mathbb{E}}|b_{N_{\ell}(nt)}|\leq C\sigma_{n}^{2}t^{\alpha^{\prime}}.$
Proof.
We start by writing
$\sum_{\ell=\ell_{n}+1}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}$ $=\int_{1/p_{\ell_{n}}}^{\infty}\sum_{j=1}^{\infty}\frac{b_{j}e^{-{nt}/{x}}}{j!}\left(\frac{nt}{x}\right)^{j}v({\rm d}x)$
$=-\int_{1/p_{\ell_{n}}}^{\infty}\sum_{j=1}^{\infty}\frac{b_{j}e^{-nt/x}}{j!}\left(\frac{nt}{x^{2}}-\frac{j}{x}\right)\left(\frac{nt}{x}\right)^{j}v(x){\rm d}x-v(p_{\ell_{n}}^{-1})\sum_{j=1}^{\infty}\frac{b_{j}e^{-ntp_{\ell_{n}}}(ntp_{\ell_{n}})^{j}}{j!}$
$=\int_{0}^{ntp_{\ell_{n}}}\sum_{j=1}^{\infty}\frac{b_{j}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z-v(p_{\ell_{n}}^{-1})\sum_{j=1}^{\infty}\frac{b_{j}e^{-ntp_{\ell_{n}}}(ntp_{\ell_{n}})^{j}}{j!}$
$=:A_{n}-B_{n}.$
Here, the first equality follows from the definition of Riemann–Stieltjes integral and also the assumption that $(p_{\ell})_{\ell\in\mathbb{N}}$ is non-increasing, the second follows from integration by parts, and the third follows from a change of variables. We claim that
(3.4) $\lim_{n\rightarrow\infty}\frac{A_{n}}{\sigma_{n}^{2}}=t^{\alpha}\int_{0}^{t\epsilon^{-{1}/{\alpha}}}\sum_{j=1}^{\infty}\frac{b_{j}e^{-z}(j-z)z^{j-1-\alpha}}{j!}{\rm d}z,$
and
(3.5) $\lim_{n\rightarrow\infty}\frac{B_{n}}{\sigma_{n}^{2}}=\epsilon\sum_{j=1}^{\infty}\frac{b_{j}e^{-t\epsilon^{-{1}/{\alpha}}}(t\epsilon^{-{1}/{\alpha}})^{j}}{j!}:=R_{\epsilon}.$
[…]
$g_{n}(z):={\bf 1}_{[0,ntp_{\ell_{n}}]}(z)\sum_{j=1}^{\infty}\frac{b_{j}e^{-z}(j-z)z^{j-1}}{j!}\frac{v(nt/z)}{v(n)},$
converges pointwise to the integrand on the right-hand side of (3.4) multiplied by $t^{\alpha}$. It remains to justify the domination condition in the dominated convergence theorem. To this end note that since
$ntp_{\ell_{n}}\sim nt\mathsf{C}_{0}\ell_{n}^{-1/\alpha}\sim nt\mathsf{C}_{0}\left(\epsilon\mathsf{C}_{0}^{\alpha}n^{\alpha}\right)^{-1/\alpha}\rightarrow t\epsilon^{-1/\alpha},$
then for some $\delta^{\prime}>0$ for $n$ large enough such that $ntp_{\ell_{n}}<t\epsilon^{-1/\alpha}+\delta^{\prime}$, and an application of Potter’s theorem (see Theorem 1.5.3 in [8]) on $v(nt/z)/v(nt)$ we have for $\delta\in(0,1-\alpha)$ and some constant $C>0$ depending on $\delta$,
$|g_{n}(z)|\leq C{\bf 1}_{[0,t\epsilon^{-{1}/{\alpha}}+\delta^{\prime}]}(z)\sum_{j=1}^{\infty}\frac{|b_{j}|e^{-z}(j+z)z^{j-1-\alpha}}{j!}\max\left\{z^{\delta},z^{-\delta}\right\}:=g(z),$
for all $n$ large enough, and $g$ is integrable.
To prove (3.5), we write
$\frac{B_{n}}{\sigma_{n}^{2}}=\frac{v(p_{\ell_{n}}^{-1})}{\sigma_{n}^{2}}\sum_{j=1}^{\infty}\frac{b_{j}(ntp_{\ell_{n}})^{j}e^{-ntp_{\ell_{n}}}}{j!},$
and observe that $ntp_{\ell_{n}}\rightarrow t\epsilon^{-1/\alpha}$ and
$v(p_{\ell_{n}}^{-1})\sim\ell_{n}.$
Then, (3.5) follows again by the dominated convergence theorem.
Next, write $\epsilon_{0}=t\epsilon^{-{1}/{\alpha}}$. Note that the right-hand side of (3.4) can be written as
$t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}(j\gamma(j-\alpha,\epsilon_{0})-\gamma(j+1-\alpha,\epsilon_{0}))}{j!}$ $=t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}(\alpha\gamma(j-\alpha,\epsilon_{0})+\epsilon_{0}^{j-\alpha}e^{-\epsilon_{0}})}{j!}$
$=t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\alpha\gamma(j-\alpha,t\epsilon^{-{1}/{\alpha}})}{j!}+R_{\epsilon},$
where we used $s\gamma(s,x)-\gamma(s+1,x)=x^{s}e^{-x}$ in the first equality. Then combining (3.4) and (3.5) we have
$\lim_{n\rightarrow\infty}\frac{1}{\sigma_{n}^{2}}\sum_{\ell=\ell_{n}+1}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}$ $=\lim_{n\rightarrow\infty}\left(\frac{A_{n}}{\sigma_{n}^{2}}-\frac{B_{n}}{\sigma_{n}^{2}}\right)=t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\alpha\gamma(j-\alpha,t\epsilon^{-1/\alpha})}{j!}+R_{\epsilon}-R_{\epsilon}$
$=\alpha t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\gamma(j-\alpha,t\epsilon^{-1/\alpha})}{j!}.$
For the second part, we have
$\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}|b_{N_{\ell}(nt)}|$ $\leq C\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}N_{\ell}(nt)^{\alpha^{\prime}}\leq C\sum_{\ell=1}^{\ell_{n}}({\mathbb{E}}N_{\ell}(nt))^{\alpha^{\prime}}=C\sum_{\ell=1}^{\ell_{n}}(p_{\ell}nt)^{\alpha^{\prime}}$
$\leq C(nt)^{\alpha^{\prime}}\ell_{n}^{1-\alpha^{\prime}/\alpha}\leq C\epsilon^{1-\alpha^{\prime}/\alpha}\sigma_{n}^{2}t^{\alpha^{\prime}},$
where in the second inequality we used ${\mathbb{E}}X^{\alpha^{\prime}}\leq({\mathbb{E}}X)^{\alpha^{\prime}}$ with $\alpha^{\prime}<1$ for non-negative random variable $X$. ∎
Proof of Lemma 3.3.
By the Cramér–Wold device it suffices to show that, for any $d\in\mathbb{N}$, distinct $t_{1}\dots,t_{d}\in[0,1]$ and any $c_{1},\dots,c_{d}\in\mathbb{R}$,
$\sum_{i=1}^{d}c_{i}\frac{\widetilde{W}_{n}^{(\epsilon)}(t_{i})}{\sigma_{n}}\Rightarrow\sum_{i=1}^{d}c_{i}\mathcal{Z}_{\alpha}^{(\epsilon)}(t_{i}).$
Note that
$\sum_{i=1}^{d}c_{i}\widetilde{W}_{n}^{(\epsilon)}(t_{i})=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-{\mathbb{E}}a_{N_{\ell}(nt_{i})}\right)$
is a summation of independent random variables, and hence we can apply the central limit theorem for an infinite triangular array of independent random variables satisfying a Lyapunov-type condition (3.7) below ([10, Theorem 8.4.1]).
We first show the convergence of variance. Since
$\mathop{\rm Var}\left(\sum_{i=1}^{d}c_{i}\widetilde{W}^{(\epsilon)}_{n}(t_{i})\right)$ $=\sum_{i,j=1}^{d}c_{i}c_{j}\mathop{\rm Cov}\left(\widetilde{W}^{(\epsilon)}_{n}(t_{i}),\widetilde{W}^{(\epsilon)}_{n}(t_{j})\right),$
it suffices to show, for $s<t$,
(3.6) $\lim_{n\to\infty}\frac{1}{\sigma_{n}^{2}}{\rm{Cov}}\left(\widetilde{W}_{n}^{(\epsilon)}(s),\widetilde{W}_{n}^{(\epsilon)}(t)\right)={\rm{Cov}}\left(\mathcal{Z}_{\alpha}^{(\epsilon)}(s),\mathcal{Z}_{\alpha}^{(\epsilon)}(t)\right).$
First, we provide the expression of the covariance function in the limit:
${\rm{Cov}}\left(\mathcal{Z}^{(\epsilon)}_{\alpha}(s),\mathcal{Z}^{(\epsilon)}_{\alpha}(t)\right)\\ =\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\gamma(k-\alpha,t\epsilon^{-1/\alpha})}{k!}-\alpha(s+t)^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(2)}(k)\gamma(k-\alpha,(t+s)\epsilon^{-1/\alpha})}{k!}.$
The calculation is essentially the same as in Proposition 2.1, replacing $\int_{0}^{\infty}$ by $\int_{0}^{\epsilon^{-1/\alpha}}$. We omit the details. Next, we compute the limit covariance. Write
${\rm{Cov}}\left(\widetilde{W}_{n}^{(\epsilon)}(s),\widetilde{W}_{n}^{(\epsilon)}(t)\right)=\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)-\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}a_{N_{\ell}(ns)}{\mathbb{E}}a_{N_{\ell}(nt)}.$
We have
$\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)$ $=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{j=1}^{\infty}\sum_{j^{\prime}=0}^{\infty}a_{j}a_{j+j^{\prime}}\frac{(nsp_{\ell})^{j}(n(t-s)p_{\ell})^{j^{\prime}}}{j!j^{\prime}!}e^{-ntp_{\ell}}$
$=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{k=1}^{\infty}\sum_{j=1}^{k}a_{j}a_{k}\frac{(nsp_{\ell})^{j}(n(t-s)p_{\ell})^{k-j}}{j!(k-j)!}e^{-ntp_{\ell}}$
$=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)(ntp_{\ell})^{k}}{k!}e^{-ntp_{\ell}}=\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}F_{\vec{a},s,t}^{(1)}(N_{\ell}(nt)).$
Thus, by Lemma 3.5 with $b_{k}=F_{\vec{a},s,t}^{(1)}(k)$, we have
$\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)\sim\sigma_{n}^{2}\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\gamma(k-\alpha,t\epsilon^{-1/\alpha})}{k!}.$
Similarly
$\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}a_{N_{\ell}(ns)}{\mathbb{E}}a_{N_{\ell}(nt)}$ $=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{j=1}^{\infty}\sum_{j^{\prime}=1}^{\infty}a_{j}a_{j^{\prime}}\frac{(nsp_{\ell})^{j}(ntp_{\ell})^{j^{\prime}}}{j!j^{\prime}!}e^{-n(s+t)p_{\ell}}$
$=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{k=1}^{\infty}F_{\vec{a},s,t}^{(2)}(k)\frac{(n(s+t)p_{\ell})^{k}}{k!}e^{-n(s+t)p_{\ell}}$
$\sim\sigma_{n}^{2}\alpha(s+t)^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(2)}(k)\gamma(k-\alpha,(t+s)\epsilon^{-1/\alpha})}{k!}.$
(Recall that $|F_{\vec{a},s,t}^{(i)}(k)|\leq Ck^{2\beta},i=1,2$ by assumption.) We have thus proved (3.6).
To complete the proof of the central limit theorem it remains to verify
(3.7) $\lim_{n\to\infty}\frac{1}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left|\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right)\right|^{2+\delta}=0,$
for some $\delta>0$. We have
$\frac{1}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left|\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right)\right|^{2+\delta}$ $\leq\frac{C}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{i=1}^{d}{\mathbb{E}}\left|a_{N_{\ell}(nt_{i})}-\mathbb{E}a_{N_{\ell}(nt_{i})}\right|^{2+\delta}$
$\leq\frac{C}{\sigma_{n}^{2+\delta}}\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{i=1}^{d}{\mathbb{E}}|a_{N_{\ell}(nt_{i})}|^{2+\delta}\leq\frac{C\sigma_{n}^{2}}{\sigma_{n}^{2+\delta}}\to 0,$
where in the second inequality, we used the fact that, by convexity and Jensen’s inequality, for random variable $X$,
(3.8) $\mathbb{E}|X-{\mathbb{E}}X|^{2+\delta}\leq 2^{1+\delta}(\mathbb{E}|X|^{2+\delta}+|\mathbb{E}X|^{2+\delta})\leq 2^{2+\delta}\mathbb{E}|X|^{2+\delta},$
and in the third inequality we used Lemma 3.5 under the assumption $b_{j}=|a_{j}|^{2+\delta}\leq Cj^{\alpha^{\prime}}$ for some $\alpha^{\prime}<\alpha$: for this it suffices to take $\delta>0$ such that $\alpha^{\prime}=\beta(2+\delta)<\alpha$. Therefore, (3.7) is verified.∎
Proof of Lemma 3.4.
We shall prove, there exist constants $C>0$ and $(d_{\epsilon})_{\epsilon>0}$ with $\lim_{\epsilon\downarrow 0}d_{\epsilon}=0$ such that
$\mathop{\rm Var}(\widetilde{T}_{n}^{(\epsilon)}(t))\leq C\sigma_{n}^{2}d_{\epsilon},\text{ for all }n\in\mathbb{N}.$
This then implies the desired result. This is then an immediate consequence of (3.3), which yields
$\mathop{\rm Var}\left(\widetilde{T}_{n}^{(\epsilon)}(t)\right)\leq\sum_{\ell=1}^{\ell_{n}}\mathbb{E}(a_{N_{\ell}(nt)}^{2})\leq C\epsilon^{1-{2\beta}/{\alpha}}n^{\alpha}t^{2\beta}.$
3.2. Tightness for the Poissonized model
To prove tightness, we proceed by another approximation of $\widetilde{W}_{n}$. Set
$j_{n}:=\left\lfloor\frac{n}{\log n}\right\rfloor,n\geq 2,$
and $j_{1}:=1$. Consider the following truncated process
$\widehat{W}_{n}(t):=\sum_{\ell=1}^{\infty}\left(\widehat{a}_{n,\ell}(t)-{\mathbb{E}}\widehat{a}_{n,\ell}(t)\right)\quad\mbox{ with }\quad\widehat{a}_{n,\ell}(t):=a_{N_{\ell}(nt)}{{\bf 1}}_{\left\{N_{\ell}(n)\leq j_{n}\right\}},t\in[0,1].$
Therefore, in order to show that $\sigma_{n}^{-1}(\widetilde{W}_{n}(t))_{t\in[0,1]}$ is tight in $D([0,1])$, we shall establish that both
$\left(\frac{\widehat{W}_{n}(t)}{\sigma_{n}}\right)_{t\in[0,1]}\quad\mbox{ and }\quad\left(\frac{\widetilde{W}_{n}(t)-\widehat{W}_{n}(t)}{\sigma_{n}}\right)_{t\in[0,1]}$
are tight. These are established in Propositions 3.6 and 3.10 below respectively.
Proposition 3.6.
Assume $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ with $\beta\in[0,\alpha/2)$.
  1. (i)
    For all $p\in[1,\alpha/{2\beta})$, we have, for some constant $C$,
    (3.9) ${\mathbb{E}}\left|\widehat{W}_{n}(t)-\widehat{W}_{n}(s)\right|^{2p}\leq C\left(\sigma_{n}^{2}|t-s|^{\alpha}+\sigma_{n}^{2p}|t-s|^{\alpha p}\right),\mbox{ for all }0\leq s<t\leq 1,n\in\mathbb{N}.$
  2. (ii)
    For some constant $C>0$, almost surely,
    (3.10) $\left|\widehat{W}_{n}(t)-\widehat{W}_{n}(s)\right|\leq C(N(nt)-N(ns)+n(t-s)),\mbox{ for all }0\leq s\leq t\leq 1.$
As a consequence, if in addition
$\beta\in\bigg{[}0,\frac{\alpha^{2}}{2}\bigg{)}\quad\mbox{ and }\quad p\in\left(\frac{1}{\alpha},\frac{\alpha}{2\beta}\right),$
the process $\sigma_{n}^{-1}\widehat{W}_{n}$ is tight in $D([0,1])$.
Proof.
We first prove (3.9). Write
$\overline{a}_{n,\ell}(t)=\widehat{a}_{n,\ell}(t)-{\mathbb{E}}\widehat{a}_{n,\ell}(t).$
By Rosenthal’s inequality (see [29]), we have
(3.11) ${\mathbb{E}}\left|\widehat{W}_{n}(t)-\widehat{W}_{n}(s)\right|^{2p}\leq C\left(\sum_{\ell=1}^{\infty}{\mathbb{E}}\left|\overline{a}_{n,\ell}(t)-\overline{a}_{n,\ell}(s)\right|^{2p}+\left(\sum_{\ell=1}^{\infty}{\mathbb{E}}\left|\overline{a}_{n,\ell}(t)-\overline{a}_{n,\ell}(s)\right|^{2}\right)^{p}\right).$
Recall (3.8). Observe also that
$\left|\widehat{a}_{n,\ell}(t)-\widehat{a}_{n,\ell}(s)\right|\leq C\left(N_{\ell}(nt)-N_{\ell}(ns)\right)^{\beta}{{\bf 1}}_{\left\{N_{\ell}(n)\leq j_{n}\right\}}\stackrel{{\scriptstyle d}}{{=}}CN_{\ell}^{\beta}(n(t-s)){{\bf 1}}_{\left\{N_{\ell}(n)\leq j_{n}\right\}},$
where the first step is in the almost sure sense and we used the property $|a_{i+j}-a_{i}|\leq Cj^{\beta}$, and the second step we used the exchangeability of Poisson random measure. Applying (3.8) first and combining all the above, we arrive at
(3.12) ${\mathbb{E}}\left|\overline{a}_{n,\ell}(t)-\overline{a}_{n,\ell}(s)\right|^{2p}\leq C{\mathbb{E}}\left|\widehat{a}_{n,\ell}(t)-\widehat{a}_{n,\ell}(s)\right|^{2p}\leq C{\mathbb{E}}\left(N_{\ell}^{2\beta p}(n(t-s)){{\bf 1}}_{\left\{N_{\ell}(n)\leq j_{n}\right\}}\right).$
Taking $p$ such that $2\beta p<\alpha$, and applying Lemma 3.7 below, we complete the proof of (3.9).
We next prove (3.10). Recall we assume that $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ for some constant $C$ for all $i,j\in\mathbb{N}$. By definition,
$\left|\widehat{W}_{n}(t)-\widehat{W}_{n}(s)\right|$ $\leq\sum_{\ell=1}^{\infty}\left|a_{N_{\ell}(nt)}-a_{N_{\ell}(ns)}\right|+\sum_{\ell=1}^{\infty}{\mathbb{E}}\left|a_{N_{\ell}(nt)}-a_{N_{\ell}(ns)}\right|$
$\leq C\sum_{\ell=1}^{\infty}\left(N_{\ell}(nt)-N_{\ell}(ns)\right)^{\beta}+C\sum_{\ell=1}^{\infty}\left({\mathbb{E}}N_{\ell}(nt)-{\mathbb{E}}N_{\ell}(ns)\right)^{\beta}$
$\leq C(N(nt)-N(ns)+n(t-s)).$
The claimed tightness now follows from Lemma 3.9 below. ∎
Lemma 3.7.
For every $r\in(0,\alpha/\beta)$, there exists $C>0$ such that
$\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}(nt)^{\beta r}{{\bf 1}}_{\left\{N_{\ell}(nt)\leq n/\log n\right\}}\right)\leq C\sigma_{n}^{2}t^{\alpha}\quad\mbox{ for all }n\in\mathbb{N},t\in[0,1].$
The upper bound is to be compared with (3.2) and (3.3) where the upper bound for the above without the indicator function was obtained. Without the indicator function, the upper bound in (3.2) is of the same order while in (3.3) we have obtained an upper bound of order $C\sigma_{n}^{2}t^{\beta r}$. The gain here is we replace $t^{\beta r}$ by $t^{\alpha}$ (recall $r\beta<\alpha$).
Proof.
We start by writing
$\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}(nt)^{\beta r}{{\bf 1}}_{\left\{N_{\ell}(nt)\leq n/\log n\right\}}\right)$ $=\int_{0}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}e^{-{nt}/{x}}}{j!}\left(\frac{nt}{x}\right)^{j}v({\rm d}x)$
$=\int_{0}^{\infty}\sum_{j=1}^{{j_{n}}}\frac{j^{\beta r}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z,$
where in the second step we used integration by parts and the substitution $z=nt/x$. Modifying the argument for the asymptotics of $A_{n}$ in (3.4) in the proof of Lemma 3.5, we can show
$\int_{0}^{1}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z\leq C\sigma_{n}^{2}t^{\alpha}\quad\mbox{ for all }n\in\mathbb{N},t\in[0,1].$
(Indeed, the left-hand side above takes the same form of $A_{n}$ with $b_{j}$ replaced by $j^{\beta r}$ and the domain of integral $[0,ntp_{\ell_{n}}]$ replaced by $[0,1]$. Recall that $\lim_{n\to\infty}ntp_{\ell_{n}}=t\epsilon^{-1/\alpha}$.) For the integral over $[1,\infty)$, we first mention an identity: for any sequence $(D_{j})_{j\geq 0}$ with $D_{0}=0$ and $j_{n}\in\mathbb{N}$,
(3.13) $\sum_{j=1}^{j_{n}}\frac{D_{j}}{j!}(j-z)z^{j-1}=\sum_{j=0}^{j_{n}}\frac{D_{j+1}-D_{j}}{j!}z^{j}-\frac{D_{j_{n}+1}}{j_{n}!}z^{j_{n}}.$
Then, it follows that for $\delta\in(0,\alpha-r\beta)$ and constant $C$ depending on $\delta$,
$\int_{1}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}e^{-z}(j-z)z^{j-1}}{j!}v\left(\frac{nt}{z}\right){\rm d}z$ $\leq v(nt)\int_{1}^{\infty}\sum_{j=0}^{j_{n}}{\frac{(j+1)^{\beta r}-j^{\beta r}}{j!}z^{j}}e^{-z}\frac{v(nt/z)}{v(nt)}{\rm d}z$
$\leq Cv(nt)\int_{1}^{\infty}\sum_{j=0}^{j_{n}}\frac{(j+1)^{\beta r}-j^{\beta r}}{j!}z^{j-\alpha+\delta}e^{-z}{\rm d}z,$
where with $D_{j}=j^{\beta r}$ we applied (3.13) in the first step (and removed the last negative term) and then applied Potter’s bound in the second step (here we needed the key property that $D_{j+1}-D_{j}\geq 0$). It remains to show that the integral above is bounded by a constant that does not depend on $n$. Indeed,
$\int_{1}^{\infty}$ $\sum_{j=0}^{j_{n}}\frac{(j+1)^{\beta r}-j^{\beta r}}{j!}z^{j-\alpha+\delta}e^{-z}{\rm d}z$
$=\int_{1}^{\infty}\sum_{j=1}^{j_{n}}\frac{j^{\beta r}(j-z)e^{-z}z^{j-1-\alpha+\delta}}{j!}{\rm d}z+\frac{(j_{n}+1)^{\beta r}}{j_{n}!}\int_{1}^{\infty}z^{j_{n}-\alpha+\delta}e^{-z}{\rm d}z$
$\leq\sum_{j=1}^{\infty}\frac{j^{\beta r}}{j!}(\alpha-\delta)\Gamma(j-\alpha+\delta)+\frac{(j_{n}+1)^{\beta r}\Gamma(j_{n}+1-\alpha+\delta)}{\Gamma(j_{n}+1)},$
which is finite since $\beta r+\delta<\alpha$. (Note that in the last step above for the first term, we first exchanged the integral and summation, next bounded $\sum_{j=1}^{j_{n}}\int_{1}^{\infty}$ by $\sum_{j=1}^{\infty}\int_{0}^{\infty}$, and then applied $j\Gamma(j-\alpha+\delta)-\Gamma(j+1-\alpha+\delta)=(\alpha-\delta)\Gamma(j-\alpha+\delta)$; if we first extend the summation and integral and then exchange the order, then Fubini’s theorem does not apply.) ∎
Remark 3.8.
The discussions around (3.11) and (3.12) are under the assumption $\beta<\alpha^{2}/2$, which is needed to establish the moment inequality on the increments in (3.9). Another method for establishing tightness is found in [12], but their approach also requires the moment estimate (3.9).
Lemma 3.9.
Let ${\boldsymbol{G}}_{n}=(G_{n}(t))_{t\in[0,1]},n\in\mathbb{N}$ be a sequence of stochastic processes defined on a common probability space with a standard Poisson process $(N(t))_{t\geq 0}$, satisfying the following two conditions.
  1. (i)
    For some $p>1/\alpha$ we have
    (3.14) ${\mathbb{E}}\left|G_{n}(t)-G_{n}(s)\right|^{2p}\leq C\left(|t-s|^{\alpha p}\sigma_{n}^{2p}+|t-s|^{\alpha}\sigma_{n}^{2}\right),\quad\mbox{ for all }s,t\in[0,1],n\in\mathbb{N}.$
  2. (ii)
    There exists a constant $C_{0}>0$ such that almost surely,
    (3.15) $\left|G_{n}(t)-G_{n}(s)\right|\leq C_{0}\left(N(nt)-N(ns)+n(t-s)\right),\quad\mbox{ for all }0\leq s\leq t\leq 1.$
Then,
$\lim_{\delta\downarrow 0}\limsup_{n\to\infty}\mathbb{P}\left(\sup_{\begin{subarray}{c}s,t\in[0,1]\\ |t-s|\leq\delta\end{subarray}}|G_{n}(t)-G_{n}(s)|\geq 9\eta\sigma_{n}\right)=0\mbox{ for all }\eta>0.$
This lemma is an adaption of [17, Lemma 3.6]. Therein, the lemma assumes a single process $G(t)$ and our lemma here is a generalization in the sense that if the process $G_{n}(t)$ takes the form $G_{n}(t)=G(nt)$, then our Lemma 3.9 becomes [17, Lemma 3.6]. However, here we shall need $G_{n}(t)=\widehat{W}_{n}(t)$, and hence we need to slightly modify the tightness argument therein.
Proof.
Fix $\eta>0$ and $\delta\in(0,1)$. Set $r:=\left\lfloor 1/\delta\right\rfloor+1$, and $t_{i}:=i\delta,i=0,\dots,r-1$ and $t_{r}:=1$. By [7, Theorem 7.4], we have
(3.16) $\mathbb{P}\left(\sup_{\begin{subarray}{c}s,t\in[0,1]\\ |t-s|\leq\delta\end{subarray}}|G_{n}(t)-G_{n}(s)|\geq 9\eta\sigma_{n}\right)\leq\sum_{i=1}^{r}\mathbb{P}\left(\sup_{s\in[t_{i-1},t_{i}]}|G_{n}(s)-G_{n}(t_{i-1})|\geq 3\eta\sigma_{n}\right).$
Fix $i\in\{1,\dots,r\}$ and we analyze the probability on the right-hand side above. Set
$x_{k,\ell}\equiv x_{k,\ell}^{(i)}:=t_{i-1}+\ell\frac{\delta}{2^{k}},\quad k\in\mathbb{N}_{0},\ell=0,\dots,2^{k}.$
For each $s\in[t_{i-1},t_{i}]$ set
$s_{k}:=\max\left\{x_{k,\ell}:\ell=0,\dots,2^{k},x_{k,\ell}\leq s\right\},\quad k\in\mathbb{N},$
and $s_{0}:=t_{i-1}$. That is, we have constructed a non-decreasing sequence: $t_{i-1}=s_{0}\leq s_{1}\leq\cdots\leq s$. We also choose
(3.17) $k_{n}:=\left\lfloor\log_{2}\left(2(e-1)\frac{C_{0}n\delta}{\eta\sigma_{n}}\right)\right\rfloor+1,$
and by triangle inequality we have
$\left|G_{n}(s)-G_{n}(t_{i-1})\right|\leq\sum_{k=1}^{k_{n}}|G_{n}(s_{k})-G_{n}(s_{k-1})|+\left|G_{n}(s)-G_{n}(s_{k_{n}})\right|.$
Now, we derive a uniform upper bound for the difference above for all $s\in[t_{i-1},t_{i}]$. For the summation on the right-hand side above, we have
$\sum_{k=1}^{k_{n}}|G_{n}(s_{k})-G_{n}(s_{k-1})|\leq\sum_{k=0}^{k_{n}}\max_{\ell=1,\dots,2^{k}}|G_{n}(x_{k,\ell})-G_{n}(x_{k,\ell-1})|.$
This is a compact way of combining two cases: when $s<t_{i}$ the summation on the right-hand side can be replaced by $\sum_{k=1}^{k_{n}}(\cdots)$ (this follows from the observation that $s_{k}-s_{k-1}=0$ or $\delta/2^{k}$ for $k\geq 1$), and when $s=t_{i}$ the left-hand side is equal to $|G_{n}(x_{0,0})-G_{n}(x_{0,1})|$ corresponding to $k=0$. We also have
$\left|G_{n}(s)-G_{n}(s_{k_{n}})\right|$ $\leq\max_{\ell=0,\dots,2^{k_{n}}-1}\sup_{s\in[x_{k_{n},\ell},x_{k_{n},\ell+1}]}C_{0}\left(N(ns)-N(nx_{k_{n},\ell})+n(s-x_{k_{n},\ell})\right)$
$\leq\max_{\ell=0,\dots,2^{k_{n}}-1}C_{0}\left(N\left(n(x_{k_{n},\ell}+\delta 2^{-k_{n}})\right)-N(nx_{k_{n},\ell})+n\delta 2^{-k_{n}}\right)$
$\leq\max_{\ell=0,\dots,2^{k_{n}}-1}C_{0}\left(N\left(n(x_{k_{n},\ell}+\delta 2^{-k_{n}})\right)-N(nx_{k_{n},\ell})\right)+\eta\sigma_{n},$
where in the first inequality we used the assumption (3.15), and in the third we recalled the choice of $k_{n}$ in (3.17) which implies $C_{0}n\delta 2^{-k_{n}}<\eta\sigma_{n}$. We then arrive at
$\limsup_{n\to\infty}\mathbb{P}$ $\left(\sup_{s\in[t_{i-1},t_{i}]}|G_{n}(s)-G_{n}(t_{i-1})|\geq 3\eta\sigma_{n}\right)$
(3.18) $\leq\limsup_{n\to\infty}\mathbb{P}\left(\sum_{k=0}^{k_{n}}\max_{\ell=1,\dots,2^{k}}|G_{n}(x_{k,\ell})-G_{n}(x_{k,\ell-1})|>\eta\sigma_{n}\right)$
(3.19) $\quad+\limsup_{n\to\infty}\mathbb{P}\left(\max_{\ell=0,\dots,2^{k_{n}}-1}C_{0}\left(N(n(x_{k_{n},\ell}+\delta 2^{-k_{n}}))-N(nx_{k_{n},\ell})\right)>\eta\sigma_{n}\right).$
(In [17, Proof of Lemma 3.6], the analysis of case where $s=t_{i}$ was missing, and it can be dealt with as shown here.) The expression of (3.19) is zero. Indeed, we have
$\mathbb{P}\left(\max_{\ell=0,\dots,2^{k_{n}}-1}C_{0}\left(N(n(x_{k_{n},\ell}+\delta 2^{-k_{n}}))-N(nx_{k_{n},\ell})\right)>\eta\sigma_{n}\right)$ $\leq 2^{k_{n}}\mathbb{P}\left(N(n\delta 2^{-k_{n}})>\frac{\eta\sigma_{n}}{C_{0}}\right)$
$\leq 2^{k_{n}}e^{n\delta 2^{-k_{n}}(e-1)-\eta\sigma_{n}/C_{0}}$
$\leq\frac{4C_{0}(e-1)n\delta}{\eta\sigma_{n}}e^{-C^{\prime}\eta\sigma_{n}},$
with $C^{\prime}=(1-1/(2(e-1)))/C_{0}$. In the second inequality above we used Markov’s inequality, and in the third we used
$2^{k_{n}}\leq\frac{4C_{0}(e-1)n\delta}{\eta\sigma_{n}}\quad\mbox{ and }\quad 2^{-k_{n}}\leq\frac{\eta\sigma_{n}}{2C_{0}(e-1)n\delta},$
which follows from the choice of $k_{n}$ in (3.17).
We obtain an upper bound for the expression (3.18). Setting $\eta_{k}=\eta/((k+1)(k+2)),k\in\mathbb{N}_{0}$, we have
$\mathbb{P}$ $\left(\sum_{k=0}^{k_{n}}\max_{\ell=1,\dots,2^{k}}|G_{n}(x_{k,\ell})-G_{n}(x_{k,\ell-1})|>\eta\sigma_{n}\right)$
$\leq\sum_{k=0}^{k_{n}}\mathbb{P}\left(\max_{\ell=1,\dots,2^{k}}|G_{n}(x_{k,\ell})-G_{n}(x_{k,\ell-1})|>\eta_{k}\sigma_{n}\right)$
$\leq\sum_{k=0}^{k_{n}}\sum_{\ell=1}^{2^{k}}\mathbb{P}\left(\left|G_{n}(x_{k,\ell})-G_{n}(x_{k,\ell-1})\right|>\eta_{k}\sigma_{n}\right).$
By moment estimates in (3.14), the last double summation is bounded by
$\sum_{k=0}^{k_{n}}\sum_{\ell=1}^{2^{k}}$ $\eta_{k}^{-2p}\frac{{\mathbb{E}}|G_{n}(x_{k,\ell})-G_{n}(x_{k,\ell-1})|^{2p}}{\sigma_{n}^{2p}}$
$\leq C\sum_{k=0}^{k_{n}}\sum_{\ell=1}^{2^{k}}\eta_{k}^{-2p}\left(|x_{k,\ell}-x_{k,\ell-1}|^{\alpha p}+\frac{|x_{k,\ell}-x_{k,\ell-1}|^{\alpha}}{\sigma_{n}^{2(p-1)}}\right)$
$\leq C\delta^{\alpha p}\sum_{k=0}^{\infty}\eta_{k}^{-2p}2^{k(1-\alpha p)}+C\delta^{\alpha}\sigma_{n}^{-2(p-1)}\sum_{k=0}^{k_{n}}\eta_{k}^{-2p}2^{k(1-\alpha)}.$
In the last expression, the first term is $C\delta^{\alpha p}$ (since the series is finite), and the second is bounded by
$C\delta^{\alpha}\sigma_{n}^{-2(p-1)}2^{k_{n}(1-\alpha)}\leq C\left(\frac{n}{\sigma_{n}^{2(p-1)/(1-\alpha)+1}}\right)^{1-\alpha}\leq C\left(\frac{n}{\sigma_{n}^{2/\alpha+1}}\right)^{1-\alpha}\to 0$
as $n\to\infty$, where in the second inequality we used the assumption $\alpha p>1$ (which implies $2(p-1)/(1-\alpha)>2/\alpha$). It then follows that
(3.20) $\limsup_{n\to\infty}\mathbb{P}\left(\sum_{k=0}^{k_{n}}\max_{\ell=1,\dots,2^{k}}|G_{n}(x_{k,\ell})-G_{n}(x_{k,\ell-1})|>\eta\sigma_{n}\right)\leq C\delta^{\alpha p}.$
Combining (3.16), (3.18), (3.19) and (3.20), we have obtained
$\limsup_{n\to\infty}\mathbb{P}\left(\sup_{\begin{subarray}{c}s,t\in[0,1]\\ |t-s|\leq\delta\end{subarray}}|G_{n}(t)-G_{n}(s)|\geq 9\eta\sigma_{n}\right)\leq C\left(\left\lfloor\delta^{-1}\right\rfloor+1\right)\delta^{\alpha p}.$
Again since $\alpha p>1$, this completes the proof. ∎
It remains to prove the tightness of $\sigma_{n}^{-1}(\widetilde{W}_{n}-\widehat{W}_{n})$.
Proposition 3.10.
With $\beta<\alpha/2$,
(3.21) $\sup_{t\in[0,1]}\frac{|\widetilde{W}_{n}(t)-\widehat{W}_{n}(t)|}{\sigma_{n}}\to 0\quad\mbox{ in probability as $n\to\infty$.}$
Proof.
To see this, pick $K>2$ and set
$\Omega_{n,K}:=\left\{\sum_{\ell=1}^{\infty}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\leq K\log n\quad\mbox{ and }\quad\sup_{\ell\in\mathbb{N}}N_{\ell}(n)\leq Kn\right\}.$
Note that,
(3.22) $\lim_{n\to\infty}\mathbb{P}(\Omega_{n,K})=1.$
Indeed, recalling that $N(n)=\sum_{\ell=1}^{\infty}N_{\ell}(n)$ is a Poisson random variable with parameter $n$, we have
$\mathbb{P}\left(\Omega_{n,K}^{c}\right)$ $\leq\mathbb{P}\left(\sum_{\ell=1}^{\infty}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\geq K\log n\right)+\mathbb{P}\left(\sup_{\ell\in\mathbb{N}}N_{\ell}(n)>Kn\right)$
$\leq\mathbb{P}\left(\sum_{\ell=1}^{\infty}N_{\ell}(n)>j_{n}K\log n\right)+\mathbb{P}\left(\sum_{\ell=1}^{\infty}N_{\ell}(n)>Kn\right)$
$\leq\mathbb{P}\left(N(n)>\frac{K}{2}n\right)+\mathbb{P}\left(N(n)>Kn\right)\to 0$
as $n\to\infty$.
Therefore, thanks to (3.22) it suffices to prove
$\lim_{n\to\infty}\mathbb{P}\left(\sup_{t\in[0,1]}\frac{|\widetilde{W}_{n}(t)-\widehat{W}_{n}(t)|}{\sigma_{n}}{\bf 1}_{\Omega_{n,K}}>\epsilon\right)=0.$
Observe that
$\sup_{t\in[0,1]}\left|{\widetilde{W}_{n}(t)-\widehat{W}_{n}(t)}\right|{\bf 1}_{\Omega_{n,K}}$
$\quad=\sup_{t\in[0,1]}\left|\sum_{\ell=1}^{\infty}\left(a_{N_{\ell}(nt)}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}-{\mathbb{E}}\left(a_{N_{\ell}(nt)}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\right)\right)\right|{\bf 1}_{\Omega_{n,K}}$
$\quad\leq\sup_{t\in[0,1]}\sum_{\ell=1}^{\infty}\left|a_{N_{\ell}(nt)}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\right|{\bf 1}_{\Omega_{n,K}}+\sup_{t\in[0,1]}\sum_{\ell=1}^{\infty}{\mathbb{E}}\left|a_{N_{\ell}(nt)}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\right|.$
For the first supremum, we have
$\sup_{t\in[0,1]}\sum_{\ell=1}^{\infty}\left|a_{N_{\ell}(nt)}{{\bf 1}}_{\left\{N_{\ell}(nt)>j_{n}\right\}}\right|{\bf 1}_{\Omega_{n,K}}$ $\leq C\sum_{\ell=1}^{\infty}N_{\ell}^{\beta}(n){{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}{\bf 1}_{\Omega_{n,K}}$
$\leq C\sup_{\ell\in\mathbb{N}}N_{\ell}^{\beta}(n)\sum_{\ell=1}^{\infty}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}{\bf 1}_{\Omega_{n,K}}\leq C(Kn)^{\beta}K\log n.$
For the second supremum, we have, for $\delta\in(0,1-\alpha)$,
$\sup_{t\in[0,1]}\sum_{\ell=1}^{\infty}{\mathbb{E}}\left|a_{N_{\ell}(nt)}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\right|$ $\leq C\sup_{t\in[0,1]}\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}^{\beta}(nt){{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\right)$
$=C\sum_{\ell=1}^{\infty}{\mathbb{E}}\left(N_{\ell}^{\beta}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}}\right)\leq C\left(\frac{\log n}{n}\right)^{\alpha+\delta-\beta}\sum_{\ell=1}^{\infty}{\mathbb{E}}N_{\ell}^{\alpha+\delta}(n)$
$\leq C\left(\frac{\log n}{n}\right)^{\alpha+\delta-\beta}\sum_{\ell=1}^{\infty}({\mathbb{E}}N_{\ell}(n))^{\alpha+\delta}\leq Cn^{\beta}\log^{\alpha+\delta-\beta}n,$
where in the second inequality we applied ${\mathbb{E}}(N_{\ell}^{\beta}{{\bf 1}}_{\left\{N_{\ell}(n)>j_{n}\right\}})\leq{\mathbb{E}}(N_{\ell}^{\beta}(N_{\ell}(n)/j_{n})^{\alpha+\delta-\beta})$ (notice $\alpha+\delta-\beta>0$). The stated result (3.21) now follows. ∎
4. De-Poissonization
Recall that $W_{n}(t)$ in (1.5) and $\widetilde{W}_{n}(t)$ in (3.1) can be re-written as follows
$W_{n}(t)$ $=\sum_{\ell=1}^{\infty}\left(a_{K_{\left\lfloor nt\right\rfloor,\ell}}-{\mathbb{E}}a_{K_{\left\lfloor nt\right\rfloor,\ell}}\right),$
$\widetilde{W}_{n}(t)$ $=\sum_{\ell=1}^{\infty}\left(a_{N_{\ell}(nt)}-{\mathbb{E}}a_{N_{\ell}(nt)}\right).$
The goal is to prove the convergence of $\sigma_{n}^{-1}(W_{n}(t))_{t\in[0,1]}$ and we have just proved the convergence of its Poissonized version $\sigma_{n}^{-1}\widetilde{W}_{n}$ in Theorem 3.1. In this section we complete the proof of Theorem 1.1 by showing that the difference of the two processes is negligible in an appropriate sense. First, the two processes can be coupled as follows. Let $(\tau_{k})_{k\in\mathbb{N}}$ denote the consecutive arrival times of the Poisson process $(N(t))_{t\geq 0}$ with $N(t):=\sum_{\ell=1}^{\infty}N_{\ell}(t)$ and $\tau_{0}=0$. Set
$\lambda_{n}(t):=\frac{\tau_{\left\lfloor nt\right\rfloor}}{n},t\in[0,1].$
Then, $(N_{\ell}(n\lambda_{n}(t)))_{t\geq 0,\ell\in\mathbb{N}}\stackrel{{\scriptstyle d}}{{=}}(K_{\left\lfloor nt\right\rfloor,\ell})_{t\geq 0,\ell\in\mathbb{N}}$. Without loss of generality, we assume further the equality is almost surely in this section. It follows that $\sum_{\ell=1}^{\infty}a_{K_{\left\lfloor nt\right\rfloor,\ell}}=\sum_{\ell=1}^{\infty}a_{N_{\ell}(n\lambda_{n}(t))}$ almost surely. Therefore, with
$\widetilde{q}_{\ell}(t):={\mathbb{E}}a_{N_{\ell}(t)},t\geq 0,\quad\mbox{ and }\quad q_{\ell}(n):={\mathbb{E}}a_{K_{n,\ell}},n\in\mathbb{N},$
we have
$\widetilde{W}_{n}(\lambda_{n}(t))$ $=\sum_{\ell=1}^{\infty}\left(a_{N_{\ell}(n\lambda_{n}(t))}-\widetilde{q}_{\ell}(n\lambda_{n}(t))\right)=\sum_{\ell=1}^{\infty}\left(a_{K_{\left\lfloor nt\right\rfloor,\ell}}-\widetilde{q}_{\ell}(n\lambda_{n}(t))\right)$
(4.1) $=W_{n}(t)+\sum_{\ell=1}^{\infty}\left(q_{\ell}(\left\lfloor nt\right\rfloor)-\widetilde{q}_{\ell}(n\lambda_{n}(t))\right).$
It is well-known that
$\lim_{n\to\infty}\sup_{t\in[0,1]}\left|\lambda_{n}(t)-t\right|=0,\mbox{ almostsurely,}$,
and hence by standard de-Poissonization argument, see for example [7] and also [17], it follows that
$\frac{1}{\sigma_{n}}\left(\widetilde{W}_{n}(\lambda_{n}(t))\right)_{t\in[0,1]}\quad\mbox{ and }\quad\frac{1}{\sigma_{n}}\left(\widetilde{W}_{n}(t)\right)_{t\in[0,1]}$
have the same limit in $D([0,1])$, and we have proved the convergence of the latter to the desired limit in the previous section. Then, in view of (4.1), to complete the proof of Theorem 1.1 it remains to establish the following. Again, the regular variation assumption (1.3) on $(p_{j})_{j\in\mathbb{N}}$ is imposed throughout without further mention.
Proposition 4.1.
Assume $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ for all $i,j\in\mathbb{N}$ with $\beta\in[0,\alpha/2)$. We have
$\sup_{t\in[0,1]}\frac{1}{\sigma_{n}}\left|\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n}(t))-q_{\ell}(\left\lfloor nt\right\rfloor)\right)\right|\to 0\mbox{ in probability as $n\to\infty$.}$
Proof.
We first show
(4.2) $\frac{1}{\sigma_{n}}\sup_{t\in[0,n^{-(1-\epsilon)}]}\left|\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n}(t))-q_{\ell}(\left\lfloor nt\right\rfloor)\right)\right|\to 0\mbox{ in probability,}$,
for $\epsilon\in(0,\alpha/2)$. We shall show that $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}(|\widetilde{q}_{\ell}(nt)|+|q_{\ell}(nt)|+|\widetilde{q}_{\ell}(n\lambda_{n}(t))|)=o(\sigma_{n})$ in probability. To see this, we first notice
(4.3) $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}\left|\widetilde{q}_{\ell}(nt)\right|\leq C\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}{\mathbb{E}}N_{\ell}(nt)\leq C\sum_{\ell=1}^{\infty}{\mathbb{E}}N_{\ell}(n^{\epsilon})\leq C{\mathbb{E}}N(n^{\epsilon})=Cn^{\epsilon},$
and $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}|q_{\ell}(nt)|\leq Cn^{\epsilon}$ can be obtained similarly. It remains to bound $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}|\widetilde{q}_{\ell}(n\lambda_{n}(t))|$. One can first show that
$\lim_{n\to\infty}\mathbb{P}\left(\sup_{t\in[0,n^{-(1-\epsilon)}]}n\lambda_{n}(t)>2n^{\epsilon}\right)=0.$
This is because the supremum of interest above is achieved at $t=n^{-(1-\epsilon)}$ and equals $n\lambda_{n}(n^{-(1-\epsilon)})=\tau_{\lfloor n^{\epsilon}\rfloor}$, which is a Poisson random variable with parameter $\lfloor n^{\epsilon}\rfloor$. Then, to show $\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}\widetilde{q}_{\ell}(n\lambda_{n}(t))/\sigma_{n}\to 0$ in probability it suffices to show
$\sup_{t\in[0,n^{-(1-\epsilon)}]}\sum_{\ell=1}^{\infty}\widetilde{q}_{\ell}(n\lambda_{n}(t))=\sup_{t\in[0,2n^{\epsilon}]}\sum_{\ell=1}^{\infty}\widetilde{q}_{\ell}(t)=o(\sigma_{n}),$
where the last step follows from (4.3). We have thus established (4.2).
It remains to show that for all $\epsilon>0$,
(4.4) $\lim_{n\to\infty}\sup_{t\in[n^{-(1-\epsilon)},1]}\frac{1}{\sigma_{n}}\left|\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n}(t))-q_{\ell}(\left\lfloor nt\right\rfloor)\right)\right|=0.$
To better illustrate the analysis, we first deal with $t=1$ (instead of considering the supremum over $t\in[n^{-(1-\epsilon)},1]$). A crucial step is to apply a local central limit theorem controlling $|\widetilde{q}_{\ell}(n)-q_{\ell}(n)|=|{\mathbb{E}}a_{N(np_{\ell})}-{\mathbb{E}}a_{B_{n,p_{\ell}}}|$ and also $|\widetilde{q}_{\ell}(n+x)-\widetilde{q}_{\ell}(n)|=|{\mathbb{E}}a_{N((n+x)p_{\ell})}-{\mathbb{E}}a_{N(np_{\ell})}|$, where $N(\lambda)$ is a Poisson random variable with parameter $\lambda>0$ and $B_{n,p}$ is a binomial random variable with parameters $(n,p)$. This step is technical and is explained in Section 4.1.
Proof for the case $t=1$. Write $\lambda_{n}\equiv\lambda_{n}(1)=\tau_{n}/n$. Set
$A_{n}:=\left\{\left|n\lambda_{n}-n\right|\leq n^{1/2}\log n\right\}.$
As a consequence of the central limit theorem we have $\lim_{n\to\infty}\mathbb{P}(A_{n}^{c})=0$, and it suffices to show
(4.5) $\lim_{n\to\infty}\frac{1}{\sigma_{n}}\left|\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n})-q_{\ell}(n)\right)\right|{\bf 1}_{A_{n}}=0,\quad\mbox{ almost surely.}$.
We have
$\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n})-q_{\ell}(n)\right){\bf 1}_{A_{n}}=\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n})-\widetilde{q}_{\ell}(n)\right){\bf 1}_{A_{n}}+\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n)-q_{\ell}(n)\right){\bf 1}_{A_{n}}=:I_{n,1}+I_{n,2}.$
We first deal with $I_{n,1}$. Set
$d_{n}:=n^{1/2}\log n\quad\mbox{ and }\quad m_{n}:=\lfloor n^{\alpha^{\prime}}\rfloor,n\in\mathbb{N},$
for some $\alpha^{\prime}>\alpha/2$ (and satisfying another condition (4.7) below). Notice
$\left|\widetilde{q}_{\ell}(n\lambda_{n})-\widetilde{q}_{\ell}(n)\right|{\bf 1}_{A_{n}}\leq\sup_{|x|\leq d_{n}}\left|{\mathbb{E}}a_{N_{\ell}(n+x)}-{\mathbb{E}}a_{N_{\ell}(n)}\right|\mbox{ almost surely.}$.
Write $J_{m}:=\{j\in\mathbb{N}:|j-m|\leq d_{m}\}$. We have
$\sup_{|x|\leq d_{n}}$ $\left|{\mathbb{E}}\left(a_{N((n+x)p)}{{\bf 1}}_{\left\{N((n+x)p)\in J_{np}\right\}}\right)-{\mathbb{E}}\left(a_{N(np)}{{\bf 1}}_{\left\{N(np)\in J_{np}\right\}}\right)\right|$
$\leq\sum_{k\in J_{np}}|a_{k}|\sup_{|x|\leq d_{n}}\left|\mathbb{P}(N((n+x)p)=k)-\mathbb{P}(N(np)=k)\right|$
$\leq C(np)^{\beta}\log n\left(O\left(\sqrt{p}\log^{2}n\right)+O\left(\frac{\log^{3}n}{\sqrt{np}}\right)\right),$
where in the second inequality we applied a local central limit theorem in Lemma 4.4. We also have
$\sup_{|x|\leq d_{n}}{\mathbb{E}}\left|a_{N((n+x)p)}{{\bf 1}}_{\left\{N((n+x)p)\notin J_{np}\right\}}\right|+{\mathbb{E}}\left|a_{N(np)}{{\bf 1}}_{\left\{N(np)\notin J_{np}\right\}}\right|\to 0$
faster than any polynomial rate (uniformly for all $p>n^{\epsilon}$). Indeed, $|a_{j}|\leq Cj^{\beta}$ and $\mathbb{P}(N(np)\notin J_{np})\to 0$ decays at rate $e^{-C\log^{2}n}$ (by Cramér–Chernoff bounds [11]), and one can show for $n$ large enough,
$\mathbb{P}(N((n+x)p)\notin J_{np})\leq\mathbb{P}\left(|N((n+x)p)-(n+x)p|>\frac{1}{2}\sqrt{(n+x)p}\log((n+x)p)\right),$
which again tends to zero at rate $e^{-C\log^{2}n}$. So, we have shown
(4.6) $\sup_{|x|\leq d_{n}}\sum_{\ell=1}^{m_{n}}\left|{\mathbb{E}}a_{N_{\ell}(n+x)}-{\mathbb{E}}a_{N_{\ell}(n)}\right|\leq Cn^{\beta}\sum_{\ell=1}^{m_{n}}p_{\ell}^{1/2+\beta}\log^{3}n+C\sum_{\ell=1}^{m_{n}}\frac{\log^{4}n}{(np_{\ell})^{1/2-\beta}}+o(n^{\alpha/2}).$
Note that when $1/2+\beta\geq\alpha$, the first summation is of order $O(n^{\beta}\log^{4}n)=o(n^{\alpha/2})$. When, $1/2+\beta<\alpha$, the first summation is of order $n^{\alpha^{\prime}(1-(1/2+\beta)/\alpha)+\beta}$, which is of order $o(n^{\alpha/2})$ if
$\alpha^{\prime}<\frac{\alpha}{2}\frac{2\alpha-4\beta}{2\alpha-1-2\beta}.$
The second summation on the right-hand side above is of order $n^{\alpha^{\prime}(1+(1/2-\beta)/\alpha)-(1/2-\beta)}\log^{4}n$, which is of order $o(n^{\alpha/2})$ when
(4.7) $\alpha^{\prime}<\frac{\alpha}{2}\frac{2+2\alpha-4\beta}{1+2\alpha-2\beta}.$
One readily checks that this condition is the most restrictive constraint on $\alpha^{\prime}$. In summary, for all $\beta\in[0,\alpha/2)$, under (4.7) the upper bound in (4.6) is of order $o(n^{\alpha/2})$ almost surely.
Next, notice that using $|a_{i+j}-a_{i}|\leq Cj^{\beta}$,
$\sum_{\ell=m_{n}}^{\infty}\left|\widetilde{q}_{\ell}(n\lambda_{n})-\widetilde{q}_{\ell}(n)\right|{\bf 1}_{A_{n}}$ $\leq\sup_{|u-n|\leq d_{n}}\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left|a_{N_{\ell}(u)}-a_{N_{\ell}(n)}\right|$
$\leq C\sup_{|u-n|\leq d_{n}}\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(|N_{\ell}(u)-N_{\ell}(n)|^{\beta}\right)$
$\leq C\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(N(d_{n}p_{\ell})^{\beta}\right).$
Assume $\alpha^{\prime}>\alpha/2$. Then, $\lim_{n\to\infty}\sup_{\ell\geq m_{n}}d_{n}p_{\ell}=0$, and therefore
$\sum_{\ell=m_{n}}^{\infty}\left|\widetilde{q}_{\ell}(n\lambda_{n})-\widetilde{q}_{\ell}(n)\right|{\bf 1}_{A_{n}}$ $\leq C\sum_{\ell=m_{n}}^{\infty}{\mathbb{E}}\left(N(d_{n}p_{\ell})\right)=Cn^{1/2}\log n\sum_{\ell=m_{n}}^{\infty}p_{\ell}$
(4.8) $\leq Cn^{1/2+\alpha^{\prime}(1-1/\alpha)}\log n=o(n^{\alpha/2}).$
We have thus proved, by choosing $\alpha^{\prime}$ satisfying (4.7) and $\alpha^{\prime}>\alpha/2$, $|I_{n,1}|=o(n^{\alpha/2})$ almost surely.
Next, we deal with $I_{n,2}$. We shall again decompose the summation into $\sum_{\ell=1}^{m_{n}}$ and $\sum_{\ell=m_{n}+1}^{\infty}$ with $m_{n}=\lfloor n^{\alpha^{\prime}}\rfloor$. However, this time the choice of $\alpha^{\prime}$ is not necessarily the same as when dealing with $I_{n,1}$ before (but instead satisfying (4.10) and (4.12) below).
We first consider $\ell\leq m_{n}$. Recall that $K_{n,\ell}$ is a binomial random variable with parameters $n,p_{\ell}$ (denoted also by $B_{n,p_{\ell}}$). This time, since $\ell\leq m_{n}$ implies $p_{\ell}\in[n^{\alpha^{\prime}/\alpha},1/2]$ we can apply Lemma 4.4. We have
$\left|{\mathbb{E}}\left(a_{N(np)}{{\bf 1}}_{\left\{N(np)\in J_{np}\right\}}\right)-{\mathbb{E}}\left(a_{B_{n,p}}{{\bf 1}}_{\left\{B_{n,p}\in J_{np}\right\}}\right)\right|$
$\leq\sum_{k\in J_{np}}|a_{k}|\left|\mathbb{P}(N(np)=k)-\mathbb{P}(B_{n,p}=k)\right|$
$\leq C(np)^{\beta}\log n\left(O\left(p\log^{2}n\right)+O\left(\frac{\log^{4}n}{np}\right)\right).$
We also have ${\mathbb{E}}\left|a_{N(np)}{{\bf 1}}_{\left\{N(np)\notin J_{np}\right\}}\right|+{\mathbb{E}}\left|a_{B_{n,p}}{{\bf 1}}_{\left\{B_{n,p}\notin J_{np}\right\}}\right|\to 0$ faster than any polynomial rate (we have discussed the first convergence before, and the second follows from Bernstein’s inequality). Thus,
$\sum_{\ell=1}^{m_{n}}\left|{\mathbb{E}}a_{N_{\ell}(n)}-{\mathbb{E}}a_{K_{n,\ell}}\right|$ $\leq Cn^{\beta}\log^{3}n\sum_{\ell=1}^{m_{n}}p_{\ell}^{1+\beta}+Cn^{\beta-1}\log^{4}n\sum_{\ell=1}^{m_{n}}p_{\ell}^{\beta-1}+o(n^{\alpha/2})$
(4.9) $\leq Cn^{\beta}\log^{3}n+Cn^{\beta-1+\alpha^{\prime}(1+(1-\beta)/\alpha)}\log^{4}n+o(n^{\alpha/2}).$
For the upper bound to be of order $o(n^{\alpha/2})$, we need to impose
(4.10) $\alpha^{\prime}<\frac{\alpha}{2}\frac{\alpha+2-2\beta}{\alpha+1-\beta}.$
For $\ell\geq m_{n}$, we have
$\left|{\mathbb{E}}\left(a_{N(np)}{{\bf 1}}_{\left\{N(np)\leq np+d_{np}\right\}}\right)-{\mathbb{E}}\left(a_{B_{n,p}}{{\bf 1}}_{\left\{B_{n,p}\leq np+d_{np}\right\}}\right)\right|$
$\leq\max_{j\leq np+d_{np}}|a_{j}|\sum_{j=1}^{\lfloor np+d_{np}\rfloor}\left|\mathbb{P}(N(np)=j)-\mathbb{P}(B_{n,p}=j)\right|$
$\leq\max_{j\leq np+d_{np}}|a_{j}|d_{\rm TV}\left(N(np),B_{n,p}\right)\leq\begin{cases}Cnp^{2},&\mbox{ if }p\leq n^{-1},\\ Cn^{1+\beta}p^{2+\beta},&\mbox{ if }p>n^{-1}.\end{cases}$
where in the second inequality we used the total variation estimate to control the difference of the two expectations: for the total variance distance between Poisson and binomial distributions we used $d_{\rm TV}(N(np),B_{n,p})\leq 2np^{2}$ [2]. We also have $|{\mathbb{E}}(a_{N(np)}{{\bf 1}}_{\left\{N(np)>np+d_{np}\right\}})-{\mathbb{E}}(a_{B_{n,p}}{{\bf 1}}_{\left\{B_{n,p}>np+d_{np}\right\}})|\to 0$ faster than any polynomial rate. Therefore,
$\sum_{\ell=m_{n}}^{\infty}\left|\widetilde{q}_{\ell}(n)-q_{\ell}(n)\right|$ $\leq Cn^{1+\beta}\sum_{\ell=m_{n}}^{\lfloor n^{\alpha}\rfloor}\ell^{-(2+\beta)/\alpha}+Cn\sum_{\ell=\lfloor n^{\alpha}\rfloor+1}^{\infty}\ell^{-2/\alpha}+o(n^{\alpha/2})$
(4.11) $\leq Cn^{1+\beta+\alpha^{\prime}(1-(2+\beta)/\alpha)}+Cn^{\alpha-1}\leq Cn^{1+\beta+\alpha^{\prime}(1-(2+\beta)/\alpha)}+o(n^{\alpha/2}).$
One readily checks that the upper bound above is of order $o(n^{\alpha/2})$, if and only if
(4.12) $\alpha^{\prime}>\frac{\alpha}{2}\frac{2-\alpha+2\beta}{2-\alpha+\beta}.$
It remains to check that there exists $\alpha^{\prime}$ satisfying both constraints (4.10) and (4.12). This is equivalent to the condition $\beta<1-\alpha/2$. But we already imposed $\beta<\alpha/2$, which is a stronger condition. That is, we have proved that, by choosing $\alpha^{\prime}$ satisfying (4.10) and (4.12), $|I_{n,2}|=o(n^{\alpha/2})$. We have completed the proof of (4.5).
Proof of (4.4). Now we extend the above upper bounds to uniform controls in both $n$ and $t$. First, one needs to modify accordingly the definition of $A_{n}$. Recall that $\lambda_{n}(t)=\tau_{\left\lfloor nt\right\rfloor}/n$ and is piecewise constant. So we have
$\sup_{t\in[0,1]}\left|n\lambda_{n}(t)-\left\lfloor nt\right\rfloor\right|=\max_{j=1,\dots,n}\left|\tau_{j}-j\right|.$
This time, we fix $\epsilon>0$ and set
(4.13) $A_{n,\epsilon}:=\bigcap_{j=\left\lfloor n^{\epsilon}\right\rfloor}^{n}\left\{|\tau_{j}-j|\leq\sqrt{j}\log j\right\}.$
Recall also that $\tau_{j}$ is the partial sum of i.i.d. standard exponential random variables. The concentration inequality for exponential random variables [34, 9] says that $\mathbb{P}(\tau_{n}-n>t)\leq e^{-C\min\{t,t^{2}/n\}}$, and we also have $\mathbb{P}(\tau_{n}-n<-t)\leq e^{-t^{2}/(2(n+t/3))}$ by Bernstein’s inequality. Thus, we have
$\mathbb{P}(A_{n,\epsilon}^{c})\leq\sum_{k=\lfloor n^{\epsilon}\rfloor}^{n}e^{-C\log^{2}k}\to 0$
as $n\to\infty$. Thus, in order to prove (4.4) it suffices to prove
(4.14) $\lim_{n\to\infty}\sup_{t\in[n^{-(1-\epsilon)},1]}\frac{1}{\sigma_{n}}\left|\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n}(t))-q_{\ell}(\left\lfloor nt\right\rfloor)\right)\right|{\bf 1}_{A_{n,\epsilon}}=0,\mbox{ almost surely.}$.
The analysis now becomes essentially the same as in the proof of the case $t=1$, except that now parameters $m_{n}$ and $d_{n}$ are chosen to depend on $t$. Our upper bounds below are uniform in $n\in\mathbb{N},t\geq n^{-(1-\epsilon)}$. Set
$m_{n}(t):=\left\lfloor(nt)^{\alpha^{\prime}}\right\rfloor\quad\mbox{ and }\quad d_{n}(t):=\left\lfloor nt\right\rfloor^{1/2}\log\left\lfloor nt\right\rfloor,$
and write
$\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n}(t))-q_{\ell}(\left\lfloor nt\right\rfloor)\right){\bf 1}_{A_{n,\epsilon}}\\ =\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(n\lambda_{n}(t))-\widetilde{q}_{\ell}(\left\lfloor nt\right\rfloor)\right){\bf 1}_{A_{n,\epsilon}}+\sum_{\ell=1}^{\infty}\left(\widetilde{q}_{\ell}(\left\lfloor nt\right\rfloor)-q_{\ell}(\left\lfloor nt\right\rfloor)\right)=:I_{n,1}(t)+I_{n,2}(t).$
In place of the analysis of $I_{n,1}$ earlier, we now have
$\left|I_{n,1}(t)\right|\leq\sup_{|u-\left\lfloor nt\right\rfloor|\leq d_{n}(t)}\sum_{\ell=1}^{m_{n}(t)}\left|\widetilde{q}_{\ell}(u)-\widetilde{q}_{\ell}(\left\lfloor nt\right\rfloor)\right|+\sup_{|u-\left\lfloor nt\right\rfloor|\leq d_{n}(t)}\sum_{\ell=m_{n}(t)+1}^{\infty}\left|\widetilde{q}_{\ell}(u)-\widetilde{q}_{\ell}(\left\lfloor nt\right\rfloor)\right|.$
This time we obtain similar bounds as in (4.6) and (4.8) (with $n$ replaced by $\left\lfloor nt\right\rfloor$), which is almost surely of order $o(n^{\alpha/2})$ under $\alpha^{\prime}>\alpha/2$ and (4.7). In place of the analysis of $I_{n,2}$ earlier, we now have
$\left|I_{n,2}(t)\right|\leq\sum_{\ell=1}^{m_{n}(t)}\left|\widetilde{q}_{\ell}(\left\lfloor nt\right\rfloor)-q_{\ell}(\left\lfloor nt\right\rfloor)\right|+\sum_{\ell=m_{n}(t)+1}^{\infty}\left|\widetilde{q}_{\ell}(\left\lfloor nt\right\rfloor)-q_{\ell}(\left\lfloor nt\right\rfloor)\right|,$
and we can obtain similar bounds as in (4.9) and (4.11), which is of order $o((nt)^{\alpha/2})$ under (4.10) and (4.12). Combining the two upper bounds we have proved (4.14). ∎
Remark 4.2.
Proposition 4.1 is much more involved than the corresponding step in earlier examples. In [17], this step is relatively straightforward (especially for the estimate of $I_{n,2}$) as with $a_{j}=1$ or $a_{j}={{\bf 1}}_{\left\{j\rm~{}is~{}odd\right\}}$, the function $\widetilde{q}_{\ell}$ and $q_{\ell}$ have explicit formulae of which one can derive a sharp control on the difference above quickly. See Lemmas 4.2 and 4.4 therein. We also notice in passing that in [12, Step 4 in the proof of Theorem 3], the statement $\mathbb{P}(Y_{n,k}^{*}(t)=Z_{n,k}^{*}\mid\Pi(n\tau)=\left\lfloor nt\right\rfloor)=1$ is incorrect (it is correct if $Y_{n,k}^{*}$ and $Z_{n,k}^{*}(\tau)$ both are replaced by the same statistic without centering). This small overlook is easy to fix therein as they considered the simple case $a_{j_{0}}=1$ for some prescribed $j_{0}$ and $a_{j}=0$ otherwise. It is also worth pointing out that they did not apply the chaining argument for the tightness of the Poissonized process, while they still needed certain moment estimates on the increments as in our (3.9).
Remark 4.3.
A natural candidate to replace $A_{n}$ (instead of $A_{n,\epsilon}$ used in (4.13)) is
(4.15) $A_{n}:=\left\{\sup_{t\in[0,1]}n^{1/2}\left|\lambda_{n}(t)-t\right|\leq\log n\right\},n\in\mathbb{N},$
for which we still have $\lim_{n\to\infty}\mathbb{P}(A_{n}^{c})=0$, given $n^{1/2}(\lambda_{n}(t)-t)_{t\in[0,1]}\Rightarrow({\mathbb{B}}_{t})_{t\in[0,1]}$ is $D([0,1])$; consequently the sequence $(\sup_{t\in[0,1]}n^{1/2}|\lambda_{n}(t)-t|)_{n\in\mathbb{N}}$ is tight. Notice that restricted to the event $A_{n}$ defined in (4.15) we have $|n\lambda_{n}(t)-nt|\leq n^{1/2}\log n$. This control could eventually lead to a uniform control in (4.14) but only for the supremum restricted to $t\in[n^{-(1/2-\epsilon)},1]$. In particular, for $t=O(n^{-1/2})$, the upper bound control is not sharp enough.
4.1. Local central limit theorems for Poisson and binomial distributions
The following lemma provides a refined control on the local central limit theorem for Poisson and binomial distributions. For example, local central limit theorem says that $\mathbb{P}(N(n)=k)=(2\pi n)^{-1/2}\exp\left(-(k-n)^{2}/(2n)\right)+o(n^{-1/2})$ and the error term $o(n^{-1/2})$ is uniformly in $k\in\mathbb{N}$. This error control is not sharp enough for our purposes and instead we shall need the following. We write $d_{n}:=n^{1/2}\log n$ and $J_{m}:=[m-m^{1/2}\log m,m+m^{1/2}\log m]\cap\mathbb{N}$.
Lemma 4.4.
Fix $\epsilon\in(0,1)$. Then, there exists $C>0$ such that
(4.16) $\max_{k\in J_{np}}\left|\mathbb{P}(N(np)=k)-\mathbb{P}(B_{n,p}=k)\right|\leq\frac{C}{\sqrt{np}}\left(p\log^{2}n+\frac{\log^{4}n}{np}\right),$
and
(4.17) $\sup_{|x|\leq d_{n}}\max_{k\in J_{np}}\left|\mathbb{P}(N((n+x)p)=k)-\mathbb{P}(N(np)=k)\right|\leq\frac{C}{\sqrt{np}}\left(\sqrt{p}\log^{2}n+\frac{\log^{3}n}{\sqrt{np}}\right),$
for all $n\in\mathbb{N},p\in[n^{-\epsilon},1/2]$.
Note that the upper constraint $p\leq 1/2$ can be relaxed to $p\leq b<1$ for any $b$ fixed. The requirement $p\geq n^{-\epsilon}$ with $\epsilon<1$ is such that $\sup_{p\in[n^{-\epsilon},1/2]}np\to\infty$, which will be used at multiple places.
Proof.
We first prove (4.16). Recall Stirling’s formula
$n!=\left(\frac{n}{e}\right)^{n}\sqrt{2\pi n}\left(1+O(n^{-1})\right).$
We start with
$\mathbb{P}(N(np)=k)=\frac{(np)^{k}}{k!}e^{-np}=\frac{1}{\sqrt{2\pi k}}\left(1-\frac{k-np}{k}\right)^{k}e^{k-np}\left(1+O(k^{-1})\right).$
Notice that
$\left(1-\frac{k-np}{k}\right)^{k}$ $=\exp\left(k\log\left(1-\frac{k-np}{k}\right)\right)$
$=\exp\left(-(k-np)-\frac{(k-np)^{2}}{2k}-\frac{(k-np)^{3}}{3k^{2}}+O\left(\frac{(k-np)^{4}}{k^{3}}\right)\right).$
The big O term on the right-hand side above is understood as follows: $O((k-np)^{4}/k^{3})$ stands for a term depending on $n,p$ and $k$, say $r_{n,p,k}$, such that for some constant $C>0$,
(4.18) $|r_{n,p,k}|\leq\frac{C(k-np)^{4}}{k^{3}},\quad\mbox{ for all }n\in\mathbb{N},p\in[n^{-\epsilon},1/2],k\in J_{np}.$
Similar interpretations apply below. Therefore, we have shown that
$\mathbb{P}(N(np)=k)$ $=\frac{1}{\sqrt{2\pi k}}\exp\left(-\frac{(k-np)^{2}}{2k}-\frac{(k-np)^{3}}{3k^{2}}\right)\left(1+O(k^{-1})+O\left(\frac{(k-np)^{4}}{k^{3}}\right)\right)$
(4.19) $=\frac{1}{\sqrt{2\pi k}}\exp\left(-\frac{(k-np)^{2}}{2k}-\frac{(k-np)^{3}}{3k^{2}}\right)\left(1+O\left(\frac{\log^{4}(np)}{np}\right)\right).$
Next, we have
$\mathbb{P}(B_{n,p}=k)$ $=\frac{n!}{k!(n-k)!}p^{k}(1-p)^{n-k}$
$=\left(\frac{np}{k}\right)^{k}\left(\frac{n(1-p)}{n-k}\right)^{n-k}\sqrt{\frac{n}{k(n-k)}}\frac{1}{\sqrt{2\pi}}\left(1+O(k^{-1})\right).$
This time, $\sqrt{n/(n-k)}=1+O\left(k/(n-k)\right)=1+O\left(p\right)$, and
$k\log{\frac{np}{k}}$ $+(n-k)\log\frac{n(1-p)}{n-k}$
$=k\log\left(1-\frac{k-np}{k}\right)+(n-k)\log\left(1+\frac{k-np}{n-k}\right)$
$=-\frac{(k-np)^{2}}{2k}-\frac{(k-np)^{3}}{3k^{2}}+O\left(\frac{(k-np)^{4}}{k^{3}}\right)-\frac{(k-np)^{2}}{2(n-k)}+O\left(\frac{|k-np|^{3}}{3(n-k)^{2}}\right).$
Notice also that
$\exp\left(-\frac{(k-np)^{2}}{2(n-k)}\right)=1+O\left(\frac{(k-np)^{2}}{2(n-k)}\right)=1+O\left(p\log^{2}(np)\right).$
We therefore have
(4.20) $\mathbb{P}(B_{n,p}=k)\\ =\frac{1}{\sqrt{2\pi k}}\exp\left(-\frac{(k-np)^{2}}{2k}-\frac{(k-np)^{3}}{3k^{2}}\right)\left(1+O\left(p\log^{2}(np)\right)+O\left(\frac{\log^{4}(np)}{np}\right)\right).$
Combining (4.19) and (4.20), we have proved the stated result (4.16).
Next, we prove (4.17). This time we have
$\mathbb{P}(N((n+x)p)=k)$ $=\frac{1}{\sqrt{2\pi k}}\left(1-\frac{k-(n+x)p}{k}\right)^{k}e^{k-(n+x)p}\left(1+O(k^{-1})\right)$
$=\frac{1}{\sqrt{2\pi k}}\exp\left(-\frac{(k-(n+x)p)^{2}}{2k}+O\left(\frac{(k-(n+x)p)^{3}}{3k^{2}}\right)\right).$
The big O term is understood similarly as in (4.18) but in addition for all $x\in\mathbb{R}$ such that $|x|\leq d_{n}$. A key observation is that for $p$ small, a typical deviation caused by $x$, $xp$, is $O(n^{1/2}p\log n)$, which is much smaller than the typical deviation $k-np$, which is $O((np)^{1/2}\log(np))$. Now,
$\frac{(k-(n+x)p)^{2}}{2k}$ $=\frac{(k-np)^{2}}{2k}+\frac{(xp)^{2}}{2k}-\frac{(k-np)xp}{k}$
$=\frac{(k-np)^{2}}{2k}+O\left(p\log^{2}n\right)+O\left(\sqrt{p}\log^{2}n\right),$
and,
$\frac{(k-(n+x)p)^{3}}{k^{2}}=O\left(\frac{(\sqrt{np}\log n)^{3}}{(np)^{2}}\right)=O\left(\frac{\log^{3}n}{\sqrt{np}}\right).$
The stated result now follows. ∎